### v3

# Amplifying Main Memory-Based Timing Covert and Side Channels using Processing-in-Memory Operations

Abstract
The adoption of processing-in-memory (PiM) architectures has been gaining momentum because they provide high performance and low energy consumption by alleviating the data movement bottleneck. Yet, the security of such architectures has not been thoroughly explored. The adoption of PiM solutions provides a new way to directly access main memory, which malicious user applications can potentially exploit. We show that this new way to access main memory opens opportunities for high-throughput timing attack vectors that are hard-to-mitigate without significant performance overhead.
We introduce IMPACT, a set of high-throughput main memory-based timing attacks that leverage characteristics of PiM architectures to establish covert and side channels. IMPACT enables high-throughput communication and private information leakage by exploiting the shared DRAM row buffer. To achieve high throughput, IMPACT (i) eliminates expensive cache bypassing steps required by processor-centric main memory and cache-based timing covert and side-channel attacks and (ii) leverages the intrinsic parallelism of PiM operations. We showcase two applications of IMPACT. First, we build two covert-channel attack variants that run on the host CPU and leverage different PiM approaches (i.e., processing-near-memory and processing-using-memory) to gain direct and fast access to main memory and establish high-throughput communication covert channels. Second, we showcase a side-channel attack that leaks private information of concurrently running victim applications that are accelerated with PiM. Our results demonstrate that (i) our covert channels achieve 12.87 Mb/s and 14.16 Mb/s communication throughput, respectively, which is up to 4.91$\times$ and 5.41$\times$ faster than the state-of-the-art main memory-based covert channels, and (ii) our side-channel attack allows the attacker to leak secrets with a low error rate. To avoid such covert and side channels in emerging PiM systems, we propose and evaluate three defenses that fundamentally eliminate the timing channel.
11footnotetext: K. Kanellopoulos and F. N. Bostancı are co-primary authors.
1 Introduction
Data movement between the main memory and the processor is a critical bottleneck present in today’s computing systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. Hence, both industry and academia have invested significant efforts to bring computation closer to where data resides, creating a paradigm called processing-in-memory (PiM)  [13, 23, 8, 24, 25, 26, 11, 27, 19, 28, 1, 2]. PiM solutions provide significant performance improvement and energy efficiency by reducing data movement.
PiM solutions take either of two approaches: (i) processing-near-memory (PnM), where the computation happens in a computation unit placed on the same chip as the memory array [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 21, 39, 40, 41, 42], which also appears in real-world products (e.g., UPMEM [43], AxDIMM [41], AiM [44], HB-PNM [45] and FIMDRAM [46]) and (ii) processing-using-memory (PuM) where the memory array performs the computation by leveraging the analog computation capability of DRAM cells [47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62]. Prior works on PnM and PuM clearly demonstrate that these emerging architectures promise great performance and energy efficiency benefits. In addition, the existence of real products that implement such techniques [43, 41, 44, 45, 46] indicates that these techniques are likely to be used in many real systems in the near future. Therefore, it is important to investigate the security aspect of such techniques and implement countermeasures to avoid potential widespread security vulnerabilities that their deployment can cause. Unfortunately, no prior work analyzes and evaluates the security of emerging PiM architectures against timing covert and side-channel attacks.
In this work, we analyze existing PiM architectures and show that the adoption of PiM solutions creates opportunities for high-throughput main memory-based attacks that are hard-to-mitigate without significant performance overheads. This is possible due to two main reasons: (i) PiM architectures provide direct access to main memory, which is essential for their performance and energy efficiency benefits, and (ii) a large set of existing practical defense mechanisms are not applicable in these PiM architectures.
1. Direct Access to Main Memory. High-throughput main memory-based timing side- and covert-channel attacks need to access main memory with high frequency to observe or change the states of shared main memory structures (e.g., DRAM row buffers). In today’s systems, it is difficult to access main memory with high frequency due to the deep cache hierarchies that filter the memory requests of attacker processes. [^1] We provide analysis as to why prior attack primitives that bypass caches do not provide fast and reliable direct main memory access in §3.2.
The adoption of PiM techniques changes how applications and computation units access main memory: (i) due to area and thermal dissipation limitations, compute units in PnM architectures have small or no caches that allow the user space applications to bypass caches more easily (i.e., with fewer memory accesses), and (ii) PuM architectures provide the ability to offload specialized PiM instructions (e.g., bulk data copy and initialization, and bitwise/arithmetic operations) to the main memory that completely bypass the deep cache hierarchies of modern processors. As a result, PiM architectures provide user space applications with fast and reliable direct main memory access that can be exploited by attackers for high-throughput timing covert- and side-channel attacks.
2. Hard-to-Mitigate with Practical Defenses. A fundamental solution to memory-based timing attacks is to eliminate the timing channel itself. Prior defenses explore fundamental solutions such as memory partitioning and enforcing constant-time memory accesses. However, these defenses incur significantly high performance overheads (as shown in Section 6). Therefore, several prior works [63, 64, 65, 66] explore practical defenses and detection mechanisms that restrict access to timing channels with cache management-based methods, such as restricting the access to cache management instructions (e.g., clflush instruction in x86 systems [63]) and detecting timing attacks based on cache access patterns (e.g., a high number of cache misses can indicate a timing attack) [64, 65, 66]. These mechanisms are inapplicable to PiM-based attacks because these attacks completely bypass the cache hierarchy.
Based on our analysis, we introduce IMPACT, a set of high-throughput In-Memory Processing-based timing Attacks that leverage the direct and fast main memory accesses enabled by PiM architectures. Using the direct access via PiM operations, IMPACT observes and exploits the shared DRAM row buffer. We demonstrate that IMPACT achieves high communication and information leakage throughput by (i) eliminating expensive cache bypassing steps required by other main memory- and cache-based timing covert- and side-channel attacks and (ii) leveraging the intrinsic parallelism of PiM operations. We showcase this finding on two covert- and one side-channel attacks. To protect future systems against IMPACT covert and side channels, we propose three defense mechanisms that fundamentally eliminate the channel itself.
IMPACT-PnM Covert Channel Variant. We build a PnM-based covert channel that uses PiM-enabled instructions [67] to directly access the main memory. The sender bypasses the locality monitor employed by the system and offloads PiM-enabled instructions for computation directly inside the main memory. The receiver measures the time it takes to access a specific set of rows to retrieve the message encoded with row buffer hits and conflicts. IMPACT-PnM achieves 12.87 Mb/s communication throughput, which is up to 4.91$\times$ faster than the state-of-the-art main memory-based attack [68].
IMPACT-PuM Covert Channel Variant. We build a PuM-based covert channel that leverages the inherent parallelism in main memory using bulk copy and initialization techniques, specifically RowClone [52], to quickly transmit messages through the shared row buffer states. This way, IMPACT-PuM transmits 16-bit large messages 14$\times$ faster than IMPACT-PnM and achieves communication throughput of 14.16 Mb/s (up to 5.41$\times$ faster than the state-of-the-art [68]).
IMPACT Side-Channel Attack. We build a PnM-based IMPACT side-channel attack that exploits PiM operations to leak private information belonging to a concurrently running application based on its memory access characteristics. We showcase this side-channel attack on read mapping, one of the fundamental applications in DNA sequence analysis where preserving the privacy of the human genomic data is crucial [69, 70, 71, 72, 73]. Our side-channel attack leverages PiM operations and leaks the private characteristics of a user’s sample genome by observing the memory access patterns of the read mapping application. Our evaluation demonstrates that our side-channel attack leaks the properties of a sample genome at a throughput of 7.57 Mb/s with 96% accuracy.
Defense Mechanisms. Based on our security analysis, we discuss and evaluate potential defense mechanisms that can mitigate IMPACT attacks by eliminating the timing channel. We discuss the performance versus security trade-offs of these defenses. We demonstrate that the two best-performing defense mechanisms lead to high performance overheads of 26% and 15% on average across five data-intensive workloads.
This paper makes the following contributions:
  • To the best of our knowledge, this is the first work that analyzes and evaluates the security of emerging PiM architectures against timing covert and side-channel attacks. We introduce IMPACT, a new set of high-throughput main memory-based timing attacks that leverage PiM operations to gain fast and reliable access to main memory.
  • We showcase two IMPACT covert channel variants, IMPACT-PnM and IMPACT-PuM that exploit PnM and PuM techniques, respectively. First, IMPACT-PnM allows the sender and the receiver to exchange messages using PiM-enabled instructions. Second, IMPACT-PuM further leverages the parallelism provided by bulk copy PuM techniques to improve the communication throughput. We show that both attacks outperform the state-of-the-art main-memory-based covert-channel attacks in terms of communication throughput.
  • We showcase an IMPACT side-channel attack on read mapping by leveraging PnM instructions. We show that our side-channel attack enables the attacker to leak sensitive information on the victim’s sample genome with high throughput and a low error rate.
  • We evaluate and discuss the trade-offs of three defense mechanisms to protect PiM-enabled systems against IMPACT’s covert- and side-channel attacks.
2 Background
2.1 DRAM Organization and Operation
Organization. Fig. 1 shows the hierarchical organization of a modern DRAM-based main memory. The memory controller connects to a DRAM module over a memory channel. A module contains one or more DRAM ranks that time-share the memory channel. A rank consists of multiple DRAM chips that operate in lock-step. Each DRAM chip contains multiple DRAM banks that can be accessed independently. A DRAM bank is organized as a two-dimensional array of DRAM cells, where a row of cells is called a DRAM row. A DRAM cell consists of 1) a storage capacitor, which stores one bit of information in the form of electrical charge, and 2) an access transistor, which connects the capacitor to the row buffer through a bitline controlled by a wordline.
Figure 1: Organization of a typical modern DRAM module.
Operation. To perform a read or write operation, the memory controller first needs to open a row, i.e., copy the data of the cells in the row to the row buffer. To open a row, the memory controller issues an activation command to a bank by specifying the row address to open. After activation is complete, the memory controller issues either read and write commands to read or write a DRAM word within the activated row. Subsequent accesses to the same row can be served quickly without fetching the data to the row buffer, creating a row hit [74]. To access data from another DRAM row in the same bank, the memory controller must first close the currently open row by issuing a bank precharge command, creating a row conflict/miss.
2.2 Processing-in-Memory
Processing-in-memory (PiM) moves computation closer to where the data resides, alleviating the data movement bottleneck between memory and the processor [13, 23, 8, 24, 25, 26, 11, 27, 19, 28, 1, 2]. There are two main PiM approaches: 1) processing-near-memory (PnM) and 2) processing-using-memory (PuM). PnM place compute logic near memory arrays (e.g., processing capability in the memory controller and logic layer(s) of 3D-stacked memory [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 21, 39, 40, 41, 42]). PnM logic can execute application regions (e.g., instructions [67, 32], functions [28], application threads [20, 54]) depending on the design. PnM computation units typically employ small caches due to 1) area constraints and 2) the ineffectiveness of caches in improving the performance of a wide range of memory-intensive applications. Therefore, various systems [32, 67, 28] offload the memory-intensive application regions on the PnM units, while executing more cache-friendly application regions on the host CPU with a deeper cache hierarchy. PuM uses analog operational properties of the memory circuitry (e.g., DRAM) to enable massively parallel in-memory computation [47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62].
2.3 Main Memory Timing Attacks
Microarchitectural timing attacks exploit the observable microarchitectural state changes to propagate messages between processes and leak secrets. These attacks measure the time to complete an operation that depends on a shared system structure (e.g., accessing a memory location by going through the shared caches) to infer the state of the shared structure.
Main memory timing attacks exploit several shared structures in the main memory. One such structure is the DRAM row buffer that acts as an internal cache inside DRAM banks. An attacker can understand if a row is recently accessed (i.e., already in the row buffer) or not by measuring the time it takes to access that row. Using this observation, DRAMA [68] builds a side-channel attack to leak keystrokes in a real system. Several prior works [68, 75, 76, 77, 78] leak DRAM address mapping functions. Other main memory-based attacks exploit the memory bus [79], and memory deduplication [80, 81] and (de)compression [82, 83, 84, 85, 86, 87, 88, 89, 90] latencies.
3 A Case for PiM-based Timing Attacks
The adoption of processing-in-memory (PiM) architectures has been gaining momentum in recent years because they provide high performance and low energy consumption by reducing data movement. Therefore, it is timely and critical to analyze and evaluate the security of PiM architectures to prevent any vulnerabilities from causing widespread susceptibility as PiM architectures become widely available. Unfortunately, no prior work analyzes and evaluates the security of emerging PiM architectures against timing covert and side-channel attacks. After a thorough analysis of existing PiM architectures, we show that the adoption of PiM solutions creates opportunities for high-throughput main memory-based timing attacks for two main reasons. First, the adoption of PiM solutions provides user space applications with fast and reliable direct access to main memory. Second, timing attacks in PiM architectures can be more difficult to detect and mitigate due to commonly employed mechanisms, like cache monitoring mechanisms that track the cache miss ratio during runtime to detect anomalies [64, 65, 66], being either inapplicable or performance-inefficient in PiM architectures.
3.1 Row Buffer Timing Channel
A DRAM row buffer acts as a direct-mapped cache inside a DRAM bank and is shared across applications that can access main memory. Accessing a row that is already in the row buffer results in a significantly lower latency because the memory controller does not need to issue additional commands to precharge the open row and fetch the target row to the row buffer. Therefore, a user space application can infer whether a row has been accessed recently by measuring the time it takes to access the row. To demonstrate the latency gap between different row buffer states, we evaluate microbenchmarks that create memory accesses that result in row buffer hits and conflicts.
We observe that a row buffer conflict takes 74 CPU cycles more than a hit, which is large enough to detect. So, a user space application can exploit the row buffer as a timing covert and side channel. Our results align with the observations of prior works [68, 77, 78].
3.2 Existing Main Memory Attack Primitives
Main memory timing attacks are powerful as they can be used to gather information on all applications that share main memory. However, these covert-channel attacks have limited throughput in traditional architectures for two main reasons: (i) accessing a memory location in main memory requires expensive cache bypassing techniques due to frequently accessed data being cached to improve performance, and (ii) even when bypassing the cache hierarchy, deep cache hierarchies employed in modern processors incur an additional high lookup latency on memory requests.
There are several ways to bypass the cache hierarchy by using one or more of four attack primitives: (i) using specialized instructions to flush a cache line (e.g., clflush), (ii) using cache eviction sets (i.e., by creating many conflicting memory requests to replace the cache line with newly-fetched cache lines), (iii) memory access requests generated by (remote) direct memory access, (R)DMA, engines, and (iv) non-temporal memory hints (e.g., movnti). However, no attack primitive provides both fast and reliable communication.
Table 1: Efficiency and Effectiveness of Attack Primitives
Low Latency Effectiveness
No Cache
Lookup
No Excessive
Memory Accesses
Timing Difference
Detectability
ISA
Guarantees
Specialized
Instructions
Eviction Sets
DMA/R-DMA N/A
Non-temporal
Memory Hints
PiM Operations
We identify four properties of an efficient (i.e., low-latency) and effective attack primitive: 1) avoiding cache lookup overhead, 2) avoiding high-latency memory accesses, 3) creating an easily detectable timing difference, and 4) reliable functionality (e.g., guaranteed to work by the ISA). Table 1 shows a comprehensive comparison of these four attack primitives in terms of these properties. First, specialized cache line flush instructions have two main disadvantages that hurt the efficiency and effectiveness of attacks that use them: 1) they put the write-back (to the main memory) latency of a cache line on the critical path to access a memory location, and 2) in some modern systems, these instructions are privileged and cannot be used by user space applications. Second, eviction sets require many load requests proportional to the number of ways in the caches, which scales with modern designs. Therefore, the attacks that use them have limited throughput. While the instructions used to create the eviction sets are part of the ISA, the eviction set is not guaranteed to evict the target cache line due to the cache replacement policy, cache prefetchers, and other design choices that are not transparent to the attacker. Third, DMA and R-DMA engines incur high overheads due to deep software stacks. As a result, attacks using this primitive need to utilize states with larger timing differences to be distinguishable. An example attack using this primitive[91] showcases a covert-channel attack based on memory contention. However, the basis of this type of covert channel is not as fine-grained as row buffer-based timing differences. Therefore, the communication throughput of these covert channels is limited (in the order of Kb/s). Fourth, the implementation of non-temporal memory hints varies across systems, as system designers are not bound to a specific approach. Consequently, the ISA does not guarantee that memory accesses related to non-temporal hints will consistently bypass caches. In contrast, PiM operations provide fast and reliable direct memory access to user space applications, which can be exploited for high-throughput covert- and side-channel attacks.
3.3 High Throughput Main Memory-based Covert Channels
To demonstrate the impact of the deep cache hierarchy latencies (i.e., cache access and eviction latencies) on an attack’s throughput, we simulate a realistic system (described in §5.1), establishing a state-of-the-art covert channel [68].
Attack Scenarios. We simulate two attacks. First, we evaluate a baseline attack that relies on cache bypassing using cache eviction sets. To transmit a bit through the row buffer, the baseline attack first evicts the corresponding cache line, and then it fetches the target row into the row buffer using an additional memory request. The baseline attack accomplishes cache line eviction by issuing N memory requests to the target cache set in an N-way cache. Note that the actual eviction latency can be much higher due to a larger number of memory accesses required to bypass more complex replacement policies employed in today’s systems or due to address translation overheads. Second, we evaluate a direct memory access attack that issues only one memory request to fetch the data to the row buffer without any cache access or evictions.
For both attacks, we measure the time it takes to send a bit through the row buffer, including cache access and eviction latencies, with encoding logic-1 values as row buffer hits and logic-0 as row buffer conflicts. We calculate the communication throughput based on the average time needed to propagate bits from the sender to the receiver. We analyze the impact of (i) the LLC size and (ii) the number of LLC ways on communication throughput to show the limitations in current and future systems with different cache designs.
Impact of LLC Size on Communication Throughput. Fig. 2 shows the impact of the LLC size on (i) the throughput of baseline and direct memory access attacks and (ii) the latency of a cache line eviction. The x-axis shows the cache configuration with increasing last-level cache (LLC) sizes and a fixed number of ways (i.e., 16). To calculate the cache access latency with increasing LLC sizes, we followed the same methodology used in prior works [92].
We make two key observations. First, the communication throughput of the direct memory access attack is $11.27$ Mb/s across all LLC sizes, which is significantly greater than the throughput of up to $2.29$ Mb/s achieved by the baseline attack. Second, the baseline attack’s throughput decreases with increasing LLC sizes since the cache access and eviction become progressively more expensive. With increasing LLC size, the cache access latency increases and, thus, amplifies the cost of eviction, which requires multiple cache accesses.
Figure 2: Impact of the LLC size on the data leakage throughput of covert channels and the eviction latency.
Impact of LLC Associativity on Communication Throughput. Fig. 3 shows the impact of LLC associativity on (i) the communication throughput of baseline and direct memory access attacks and (ii) the eviction latency of a cache line. The x-axis shows the increasing number of LLC ways (from 2 to 128) while the total LLC size is fixed at 16 MB (modern processors typically employ 8 cores and a 2MB LLC slice per core). Increasing the number of LLC ways increases the number of memory requests required to evict a cache line from the cache. We observe that the baseline attack’s communication throughput decreases significantly with an increasing number of LLC ways due to the higher cache eviction latency.
Figure 3: Impact of the number of LLC ways on the data leakage throughput of covert channels and the eviction latency.
We conclude that attacks that communicate with main memory through caches are greatly limited in throughput due to expensive cache access and eviction latencies, which worsen with increasing size and associativity. However, a covert channel based on direct memory access can achieve high throughput independent from cache configurations as it does not require bypassing the cache hierarchy.
The adoption of PiM solutions can fundamentally enable high-throughput timing attacks through main memory by providing direct access to the main memory to user space applications. Therefore, it is crucial to carefully consider the security implications of PiM architectures and develop robust security measures to protect against potential attacks.
4 IMPACT: A Set of High-Throughput Main Memory Timing Attacks
This section introduces IMPACT, a set of high-throughput main memory-based timing attacks that take advantage of the direct memory access enabled by the emerging processing-in-memory paradigm. IMPACT achieves high communication and information leakage throughput by (i) eliminating expensive cache bypassing steps required by other main memory- and cache-based timing covert- and side-channel attacks and (ii) leveraging the intrinsic parallelism of PiM operations. In this work, we present two covert-channel and one side-channel attacks enabled by processing-near-memory (PnM) and processing-using-memory (PuM) solutions.
4.1 PnM-based IMPACT Covert-Channel Attack
Baseline PnM Architecture. PiM-Enabled Instructions (PEI)  [67] is a PnM architecture that expands the instruction set architecture of the host processor with PIM-enabled instructions (PEIs). PEIs can be executed either on the host processor or on the in-memory computation logic. We choose PEI as the baseline PnM architecture as it introduces a simple and effective model for PnM, and our attack can be generalized for other PnM architectures with similar design components (e.g., FIMDRAM [46]).
The baseline PEI architecture has two key components. First, it integrates a PEI Computation Unit (PCU) near each DRAM bank and inside the CPU. In-memory PCUs are shared among all the cores of the host processor and are responsible for executing PEIs near the respective DRAM bank. Second, the PEI architecture has a PEI Management Unit (PMU) that monitors data locality in application regions and decides where to map each region (either to host-side or memory-side PCUs). This way, PEI architecture can execute regions with high locality on the host side to benefit from caches. PEI adds an ignore flag per locality monitor entry to ignore the first hit in the locality monitor [93]. This is because if a locality monitor entry is allocated for a PIM operation, it is often too aggressive to consider the operation as having a high locality on its first hit. This bit allows IMPACT-PnM to easily bypass the locality monitor even with a small virtual address range at its disposal.
Attack Overview. Before the attack, the sender and the receiver co-locate their data in the same DRAM banks. To achieve this, one process uses memory massaging techniques to place its data in the same bank as the other process in main memory, as used in many prior works [76, 68, 94, 95].
IMPACT-PnM uses PEI operations and exploits the latency difference between a DRAM row buffer hit and a conflict in a shared DRAM bank to exchange messages between the sender and the receiver. The sender transmits bits of the message to the receiver in multiple batches of M bits, where the value of each bit is encoded as interference in the row buffer of a DRAM bank. In this attack, we encode logic-1 as interference in the DRAM bank (i.e., a row buffer conflict), and logic-0 as no interference (i.e., a row buffer hit).
Fig. 4 and Listing 1 show the end-to-end flow of IMPACT-PnM and all the steps involved in establishing the covert channel between the sender and the receiver.
Figure 4: Flow of the PnM-based Covert Channel Attack.
1 Sender:
2 message = [0: N-1]; //N banks and N-bit msg.
3 batch_size = M; //M-bits per batch where M<N
4 init_semaphore(0);
5 barrier(); // synchronize with receiver
6 bank=0;
7 for every batch:
8 for each bit in batch_size;
9 if message[bank] == 1:
10 pim_add(row[bank]);
11 else:
12 NOP();
13 bank++;
14 memory_fence(); // signal receiver
15 increment_semaphore(); // to start probing
16
17 Receiver:
18 init_row = initialize_DRAM_rows_with_PEIs();
19 bank = 0;
20 barrier(); // synchronize with sender
21 for every batch:
22 block_on_semaphore();
23 //if semaphore==0, wait
24 for each bit in batch_size;
25 timer(start);
26 pim_add(init_row[bank]);
27 timer(end);
28 latency = end - start;
29 if latency > threshold:
30 result = ’1’;
31 else:
32 result = ’0’;
33 bank++;
Listing 1: Pseudocode of PnM-based Covert-Channel Attack.
Step 1. The receiver issues PEIs to initialize each DRAM bank by activating a predetermined row (Line 18). The receiver either uses unused cache lines of the activated row or sets the ignore flag in the corresponding entry in order to bypass the locality monitor. Then, the sender and the receiver synchronize on a barrier (Lines 5 and 20), which ensures that they start the communication turn at the same time.
Step 2. The sender starts the communication turn by issuing PEI instructions to activate a row in each DRAM bank (Lines 7-12) or not based on the bit value that will be transmitted. To transmit a 0 ( 1), the sender does not activate any row in the corresponding DRAM bank to avoid causing interference in the row buffer (e.g., Row 0xB remains in the row buffer  2). To transmit a 1 ( 3), the sender activates a different row to cause interference in the corresponding DRAM bank (i.e., a row buffer conflict between Row 0xD and Row 0xE  4). At the end of the communication turn (after transmitting M bits), the sender first executes a memory fence (Line 14) to ensure that all the PEI instructions have been executed and then signals the receiver to start probing the DRAM banks (Line 15).
Step 3. The receiver starts probing the DRAM banks by issuing PEI instructions to activate the same row as the one activated during initialization (Line 25). Then, the receiver measures the time it takes to complete the request (Line 27). If the latency exceeds a predetermined hit threshold, the receiver senses the interference in the row buffer and decodes the bit as a 1 (Lines 28-29  5, 6). Otherwise, the access is considered a row buffer hit and the receiver decodes the bit as a 0 (Lines 30-31  7, 8). The sender and the receiver repeat the same steps for each batch until the entire message is transmitted.
Sender-Receiver Synchronization. The sender and the receiver synchronize to 1) ensure their actions do not interfere with each other and 2) overlap the latencies of their operations to increase the throughput of the attack. To do so, the sender and the receiver use a semaphore. The value of the semaphore indicates the number of batches that the sender has completed transmitting, but the receiver has not yet probed. The receiver blocks on the semaphore until the sender increments it (i.e., does not start probing). When the sender completes the transmission of a batch, they increment the semaphore to signal the receiver to start. Upon receiving the signal, the receiver decrements the semaphore and starts probing the DRAM banks. In parallel, the sender can start transmitting the next batch that utilizes a different set of DRAM banks. This way, they can overlap the transmission latency and the probing latency. When the receiver completes probing, they read the semaphore again, and the process repeats until the entire message is transmitted. In our attack implementation (§5.1), we use POSIX semaphores to implement the sender-receiver synchronization and use a batch size of 4 bits.
4.2 PuM-based IMPACT Covert-Channel Attack
Baseline PuM Architecture. We assume a PuM architecture that provides user applications with RowClone [52] that enables bulk data movement operations (e.g., bulk data copy and initialization) within DRAM’s subarray boundaries by exploiting the internal operations of DRAM. [^2] There could be near-memory RowClone variants that move data without using operational principles of DRAM as well. However, in this attack, we assume a PuM-based RowClone operation. RowClone’s Fast Parallel Mode works by issuing back-to-back two activation commands: the first activation copies the content of a source row (i.e., src) into the row buffer. The second command connects the destination row (i.e., dst) to the bitlines. Since the row buffer already stored the src’s data, by the time the dst is activated, the data in dst is overwritten by src’s data.
To execute RowClone, we assume that the user space application specifies a source and a destination virtual address range and the corresponding mask to copy only specific parts of the source to the destination. This way, the memory controller receives one RowClone request and breaks it into parallel RowClone requests, one for each set bit of the mask.
Attack Overview. PuM-based covert channel attack (IMPACT-PuM) exploits in-DRAM bulk copy operations to establish a covert channel between a sender and a receiver. The sender transmits a message by executing a RowClone operation with three operands: (i) a source address range, (ii) a destination address range, and (iii) a mask to select which parts of the ranges will be copied. Source and address ranges correspond to physical pages spanning across all DRAM banks and by using the mask, the sender can encode the transmitted message by causing contention only in selected DRAM banks. The receiver repeatedly issues RowClone operations and measures the latency of these operations. This way, if the sender issues a RowClone operation to a DRAM bank, it induces contention, and the receiver detects a higher latency for its RowClone operation, indicating the transmitted bit is logic-1. Otherwise, the receiver detects a lower latency for the RowClone operation, indicating that the corresponding bit is a logic-0.
Fig. 5 and Listing 2 show the end-to-end flow of IMPACT-PuM and all the steps involved to establish the covert channel between the sender and the receiver.
Figure 5: PuM covert channel attack flow.
Step 1. First, the receiver initializes all DRAM banks by issuing a RowClone operation (Line 16). Then, the sender and the receiver synchronize on a barrier (Lines 10 and 17) which ensures that they start the communication turn at the same time.
Step 2. The sender uses a single RowClone operation to transmit an N-bit message (where N is the number of banks) to the receiver in parallel and uses the mask of the RowClone operation to encode the message (Lines 11-12). The sender encodes the message by setting the corresponding bits in the mask to 1 or 0 ( 1, 2). To send a 1, the sender triggers RowClone in the corresponding DRAM bank and induces contention by altering the activated row in the row buffer ( 3). To send a 0, the sender does not trigger RowClone in the corresponding DRAM bank and does not induce contention ( 4). The sender synchronizes with the receiver using a barrier operation (Lines 13 and 18).
1 Memory Controller:
2 row_clone(src, dst, mask):
3 for each DRAM bank do:
4 if mask[bank] == 1 then:
5 row_clone(bank, src, dst);
6 else:
7 NOP();
8
9 Sender:
10 barrier_1();
11 mask = init(message);
12 row_clone(src, dst, mask);
13 barrier_2();
14
15 Receiver:
16 init_DRAM_rows_with_RowClone();
17 barrier_1();
18 barrier_2();
19 for each DRAM bank do:
20 mask = 1 << bank;
21 timer(start);
22 row_clone(src, dst, mask);
23 timer(end);
24 latency = end - start;
25 if latency > threshold:
26 result = ’1’;
27 else:
28 result = ’0’;
Listing 2: Pseudocode of PuM-based Covert-Channel Attack.
Step 3. The receiver in IMPACT-PuM detects contention in DRAM banks by issuing RowClone operations and measuring their latency. The receiver continuously issues one RowClone operation for each DRAM bank (in contrast to the sender who issues only a single RowClone operation), to the initial virtual address range (i.e., rows initialized in Step 1) and measures the latency of each operation ( 2, Line 24). If the sender induces contention in a DRAM bank, the receiver observes high latency for the RowClone operation which indicates that the corresponding bit in the message is set to 1 ( 3, Lines 25-26). Otherwise, the receiver detects a fast RowClone operation which indicates that the bit is set to 0 ( 4, Lines 27-28).
Advantage over IMPACT-PnM. In IMPACT-PuM, the sender can use a single RowClone operation to transmit an N-bit message to the receiver in parallel, where N is the number of DRAM banks. Thus, IMPACT-PuM requires less computational resources to transmit the same amount of information compared to IMPACT-PnM. For example, using a single thread, the sender can activate all N DRAM banks and transmit N bits of information in parallel. In contrast, in IMPACT-PnM, the sender needs to execute one PEI per bit.
4.3 PnM-based Side-Channel Attack on Read Mapping
We build a PnM-based side-channel attack that exploits PiM operations to leak private information of a concurrently running application. To demonstrate the IMPACT side-channel attack, we target a DNA sequence analysis application.
Read Mapping. DNA is one of the unique identifiers of an individual [73], such that even a small fraction of DNA can include sufficient information for linking to the owner’s identity [70, 71]. Therefore, preserving privacy while processing genomic data carries utmost importance to protect sensitive information. Sequencing the DNA of a biological sample produces short DNA fragments called reads. The reads are then processed by read mapping, which is a fundamental operation to find the matching locations of reads in the reference genome that acts as a template to construct the complete genome of the sample.
Read mapping (RM) has two key steps: (i) seeding and (ii) alignment. During the seeding step, RM identifies the possible locations of the reads on the reference genome based on sequence similarities between the reads and the reference genome. Seeding involves hashing small segments of the DNA sequences (i.e., seeds) and probing a hash table that is built from the reference genome. After seeding, RM aligns the reads with their candidate regions on the reference (i.e., potential reference segments) using dynamic programming-based methods to find each read’s best matching position among all possible locations in the alignment step.
RM requires a large amount of data movement to access different parts of the hash table and corresponding regions on the reference genome. Hence, several prior works aim to improve the performance of RM by offloading seeding and alignment to PiM-enabled systems to reduce the data movement overheads [33, 96, 97, 98, 99, 100, 101, 102]. In this work, we demonstrate that an attacker can launch a high-throughput side-channel attack against RM by utilizing PiM instructions and quickly leak a large amount of information about the sample genome.
Attack Overview. We assume that the victim is running a RM application on a PiM-enabled system and the attacker is able to run a malicious application on the same system. The attacker and the victim are using the same read mapping tool (e.g., minimap2 [103]). The read mapping tool constructs a hash table that contains information about the seed locations in the reference genome, and allows each user in the system to probe the hash table via queries during read mapping. We assume the hash table is distributed across multiple DRAM banks. This assumption is realistic as many modern DRAM address mapping schemes (e.g., [104, 105, 106, 107]) aim to interleave consequent memory chunks across different DRAM banks to exploit bank-level parallelism [108, 109].
Fig. 6 shows the high-level overview of the attack. The victim application extracts the seeds from the reads and offloads the corresponding seeding and alignment step to the PiM-enabled system  1. During the seeding step, the corresponding DRAM row gets activated (i.e., Index 0-5 of the hash table)  2. The attacker continuously uses PnM instructions to probe the shared hash table in an attempt to identify if the victim application is accessing a specific index of the hash table  3. If the attacker observes that the victim application is accessing a specific index, it retrieves partial or exact information about the potential locations of the reads in the reference genome. As shown in previous works on DNA imputation [110, 111, 112, 113], the attacker can use the leaked information in a completion attack [71] to infer properties about some regions of the private sample genome  4.
Figure 6: High-level overview of IMPACT attack model on PiM-enabled RM.
5 Evaluation
5.1 Methodology
Covert Channel Threat Model. We assume a scenario where a sender and a receiver operate on the same PiM-enabled system to exchange information. PiM instructions are available to user space applications without any restrictions. The receiver has access to cpuid and rdtscp instructions, enabling high-precision measurement of memory access latencies. For IMPACT-PnM, the sender and the receiver co-locate data in the same DRAM banks. For IMPACT-PuM, (i) the sender allocates two virtual address ranges whose corresponding physical pages span across all DRAM banks, and (ii) RowClone guarantees atomicity: when the sender issues a RowClone operation that spans across multiple banks, all individual bank-level RowClone operations need to be completed before another DRAM operation is executed.
Side Channel Threat Model. We assume a scenario where an attacker process operates on the same PiM-enabled system as a victim process and leaks information about the victim process. We use an RM implementation based on minimap2 [103], and we modify the implementation to offload the seeding and alignment step to the PiM-enabled system. We assume that the alignment step includes chaining. We use the human genome as the reference and compare it against synthetic sample genomes. We experiment with multiple seed sizes and report the best-performing selection.
Proof-of-Concept. We construct a proof-of-concept of all IMPACT variants using the Sniper Multicore Simulator [114]. [^3] There is no commercially available system that follows a similar architecture as PEI or any modern system that enables end-to-end RowClone operations. We extend Sniper to (i) accurately model all internal DRAM operations, (ii) emulate and simulate the PiM interface and PiM operations, and (iii) emulate the functionality of the cpuid and the rdtscp instructions to measure execution cycles from the user space.
Table 2 shows the configuration of our simulation setup. Before launching the attack, the sender and the receiver threads warm up by executing both attacks to avoid compulsory DRAM accesses for instructions or page table walks. For IMPACT-PnM, we assume that both the attacker and the receiver use PEIs to perform an operation that takes $\sim 3$ cycles to execute (in addition to the actual DRAM access).
Comparison Points for Covert-Channel Attacks. To showcase the effectiveness of IMPACT, we evaluate the leakage throughput of seven different covert-channel attacks.
(i) DRAMA-clflush: The row buffer-based covert-channel attack proposed in [68] which uses clflush instructions to send memory requests directly to main memory. clflush only probes the LLC to flush the cache line.
(ii) DRAMA-Eviction: The row buffer-based covert channel attack proposed in [68] which uses cache eviction sets to send memory requests directly to main memory.
(iii) Streamline: State-of-the-art cache covert channel proposed in [115] which enables a flushless communication.
(iv) DMA Engine: A row buffer-based attack that uses the DMA engine to bypass the caches. We assume a powerful attacker can infer the DRAM access time directly from the user space by ignoring context switch and OS-related latencies.
(v) PnM-OffChip: A PnM-based attack in a PnM architecture with locality monitor and a state-of-the-art off-chip predictor [116]. If the data is predicted to have high locality, the computation is performed by the CPU instead of the near-memory compute units, leading to reduced attack throughput.
(vi) IMPACT-PnM (§4.1): An attack that bypasses the locality monitor and offloads PEI instructions directly to compute units near DRAM banks in order to establish a covert channel.
(vii) IMPACT-PuM (§4.2): An attack that leverages RowClone to offload bulk-copy operations directly to main memory and establish a covert channel.
For a fair comparison against DRAMA and Streamline, we showcase the upper bound of the communication throughput achieved by each attack. To calculate their throughput, we use our simulation infrastructure to extract parameters such as the LLC hit latency, average LLC miss latency, cache lookup latency, cache hit/miss ratio, and feed them in an analytical model that calculates the maximum throughput of the attack. We validate the model against the real system throughput of DRAMA and Streamline (e.g., Streamline [115] reports a maximum communication throughput of 1.8 Mb/s and we model the upper bound as 2.7 Mb/s for the smallest LLC).
Noise Sources in Simulation Setup and Error Rate. We simulate hardware prefetchers and page table walkers to induce noise in the simulated system. We measure the throughput of each attack only based on the successfully leaked data (i.e., incorporating the error rate in our measurements).
Table 2: Simulation Configurations
CPU 4-core, 4-way, OoO x86 core, 2.6GHz
MMU L1 DTLB (4KB): 64-entry, 4-way assoc, 1-cycle access, L1 DTLB (2MB): 32-entry, 4-way assoc, 1-cycle access, L2 TLB: 1536-entry, 12-way assoc, 12-cycle access
L1 Cache I-cache: 32 KB; L1D Cache: 32 KB, 8-way assoc, 4-cycle access latency, LRU, IP-stride prefetcher [117]
L2 Cache 1 MB, 16-way assoc, 12-cycle latency, SRRIP [118], Streamer [119]
L3 Cache 2 MB/core, 16-way assoc, 32-cycle latency, SRRIP [118]
Main Memory DDR4-2400, 16 Banks, 4 Ranks, 1 Channel, Row size = 8192 bytes, $t_{RCD}$ = 13.5 ns, $t_{RP}$ = 13.5 ns , $t_{RC}$ = 13.5 ns, Open Row policy, Row Timeout = 100 ns
5.2 Validation of Proof-of-Concept
We first demonstrate the PoC results and how we successfully mount the attack in the simulated system. Fig. 7 shows the latency (in cycles) measured by the receiver after executing (a) a PEI instruction and (b) a RowClone operation, one for each of the 16 banks assuming the message is 16-bits large, in IMPACT-PnM and IMPACT-PuM, respectively. We observe that the receiver can successfully distinguish whether the sender causes contention in the bank in both IMPACT-PnM and IMPACT-PuM. The receiver can decode the complete message using a threshold value of 150 cycles as follows. If the latency is less than 150 cycles (i.e., row buffer hit), the receiver decodes a bit as logic-0, otherwise (i.e., row buffer conflict) it decodes the bit as logic-1.
Figure 7: Latency (in cycles) measured by the receiver to execute (a) a PEI instruction and (b) a RowClone instruction for each bank in order to decode a 16-bit message.
5.3 Covert-Channel Attack Throughput Analysis
Fig. 8 shows the throughput achieved by different covert-channel attacks when one sender and one receiver thread run concurrently across systems with increasing LLC sizes. We make five key observations. First, IMPACT-PnM and IMPACT-PuM achieve significantly higher throughput than all other attack vectors. IMPACT-PnM and IMPACT-PuM achieve communication throughputs of 12.87 Mb/s and 14.16 Mb/s, respectively, irrespective of the cache size. They support up to 4.91$\times$ and 5.41$\times$ of the throughput achieved by the state-of-the-art main memory-based attack, DRAMA-clflush. Second, IMPACT-PuM achieves 10% higher throughput than that of IMPACT-PnM since the sender can issue all RowClone requests in parallel, whereas in IMPACT-PnM, the sender has to send the PiM instructions in a sequential manner. Third, as the LLC size increases, the throughput of DRAMA-eviction, DRAMA-clflush, and Streamline decreases significantly. This is because the larger the LLC size, the higher the latency it takes to access the LLC. Fourth, the DMA engine attack achieves a throughput of 5.27 Mb/s irrespective of the cache size but is almost 2.4$\times$ slower than IMPACT-PnM due to its OS-related overheads (e.g., context switch overhead, OS instructions). Fifth, the PnM-OffChip attack achieves a maximum throughput of 12.64 Mb/s. As the LLC size increases, the throughput is reduced to 10.64 Mb/s because the off-chip predictor decides to cache more data when the LLC is large. We conclude that IMPACT achieves significantly higher throughput than the existing state-of-the-art main memory-, cache- and DMA engine-based covert-channel attacks.
Figure 8: Throughput across increasing LLC sizes.
Breakdown of Throughput Improvements. Fig. 9 shows the time (in CPU cycles) it takes for the sender to send the message and for the receiver to read the message in IMPACT-PnM and IMPACT-PuM. We observe that in IMPACT-PuM, the sender can send the message 14$\times$ faster than the sender in IMPACT-PuM since it can issue all RowClone requests in parallel. In IMPACT-PnM, even though the sender is significantly slower than the sender of IMPACT-PuM, the sender and the receiver overlap, at the cost of synchronization, to hide the latency of the sender. This is why IMPACT-PnM achieves only 10% lower throughput than IMPACT-PuM.
Figure 9: Time (in cycles) it takes (i) for the sender to send the message and (ii) for the receiver to read the message in IMPACT-PnM and IMPACT-PuM.
5.4 Side-Channel Attack Throughput Analysis
Fig. 10 shows the throughput and the error rate of the side-channel attack against read mapping (RM) across different numbers of DRAM banks in the left and right y-axis, respectively. We measure the throughput of the attack based on the correct guesses and calculate the error rate based on the number of incorrect guesses. The exact end-to-end accuracy of identifying the properties of the sample genome depends on the chaining mechanism and the imputation algorithm [110, 111] used by the attacker, and it is out of the scope of this work.
We make two key observations. First, with a PiM-enabled DRAM with 1024 banks, the attacker can leak information about the sample genome at a throughput of 7.57 Mb/s and an error rate of $<5$%. Second, as the number of banks increases to 8196, the throughput of the attack decreases to 2.56 Mb/s and the error rate increases to $<15$%. This is because the attacker has to probe more banks, which makes the attack more prone to noise and reduces its bandwidth.
However, at the same time, the leaked information provides more exact information about the sample genome since the number of hash table entries per bank decreases, reducing the number of potential locations of the reads in the reference genome per bank. For example, in a PiM-enabled DRAM device with 1024 banks, the victim activates a row with 16 hash table entries. Assuming the attacker identifies the activated bank correctly, they determine that one of the 16 hash table entries was targeted by the victim. In contrast, in a PiM-enabled device with 2048 banks, the victim activates a row with 8 hash table entries. When the attacker identifies the activated bank correctly, they determine that one of the 8 hash table entries was targeted by the victim, which leaks more precise information on the sample genome.
Figure 10: Leakage throughput and error rate of the RM side-channel attack across different numbers of DRAM banks.
6 Defenses against IMPACT
This section discusses the effectiveness of three defense mechanisms that eliminates the row buffer-based timing channel.
Memory Partitioning (MPR). Bank-level memory partitioning enforces that each DRAM bank can only be accessed by one process that it is allocated to. In IMPACT, this limits the sender’s ability to transmit information across different DRAM banks and the receiver’s ability to observe contention created by the sender. However, this approach has three major drawbacks. First, it can lead to limiting the number of concurrently running applications due to the fixed number of DRAM banks available. Second, it can cause memory underutilization due to allocating DRAM bank-sized memory to applications. Third, it disables data sharing, which can significantly increase data duplication and memory underutilization, and lead to frequent swapping from the disk when the memory space is full.
Closed Row Policy (CRP). Enforcing a memory controller policy that closes a DRAM row after each operation can be an effective defense against IMPACT. By closing the DRAM row after each operation, the memory controller ensures that subsequent memory accesses cannot access the same row without requiring a row activation. Closed row policy disrupts the timing patterns exploited by the attacker in IMPACT. However, it also leads to significant performance overheads, even for workloads with low data locality, as every DRAM access leads to a row buffer miss.
Constant-Time DRAM Accesses (CTD). Another defense mechanism is to enforce a constant-time DRAM access (CTD) policy. When a memory access is issued, the memory controller waits for a fixed amount of time before returning the data to the processor. This fixed amount of time is equal to the worst-case DRAM access latency. By enforcing CTD, the memory controller ensures that all memory accesses take the same amount of time, regardless of the memory location. Thus, obfuscates the timing patterns exploited by the attacker in IMPACT. However, this leads to significant performance overheads as every DRAM access has the worst-case latency.
Performance Evaluation. We evaluate the performance of CRP and CTD compared to the baseline open-row policy for 5 multi-programmed graph applications which (i) share the input graph (i.e., access to the same DRAM bank) and (ii) experience a mix of low and high data locality (i.e., the number of last-level cache misses per kilo instructions). We use a 2-core system where each core runs a different instance of the same application (e.g., core 0 runs an instance of PageRank and core 1 the same instance of PageRank using the same input). We evaluate five workloads from GraphBIG benchmark suite [120]: Betweenness Centrality (BC; Data set size: 7.8 GB; LLC MPKI: 0.57), Breadth-First search (BFS; 7.6 GB; 38.59), Connected components (CC; 7.9 GB; 45.2), Triangle counting (TC; 8.2 GB; 5.08), PageRank (PR; 7.0 GB; 1.86).
Fig. 11 shows the performance overhead of CRP and CTD compared to the open-row policy. We make two key observations. First, CTD leads to the highest performance degradation, 26% on average, compared to the open row policy. Second, CRP leads to an average performance degradation of 15% compared to the open row policy but leads to low performance overheads in workloads that do not benefit from the open row policy such as TC, CC and BFS. We conclude that fundamental defense mechanisms that eliminate the timing channel lead to high performance overheads, and there is a dire need for high-performance defense mechanisms.
Figure 11: Performance overhead of Closed-Row policy and Constant-Time policy compared to Open-Row policy.
7 Related Work
To the best of our knowledge, this is the first work that evaluates the timing attack vulnerabilities in PiM systems.
Security Implications of PiM Systems. A set of prior works [121, 122, 123] discuss specific security vulnerabilities in PiM systems. Arafin et al. [121] survey and theoretically analyze security vulnerabilities, focusing on five specific attacks. In this work, we quantitatively and qualitatively analyze and evaluate main memory-based covert and side channels in detail. Wang et al. [123] extracts neural network architectural information from an RRAM accelerator using power side channels. Ensan et al. [122] proposes an attack on RRAM-based PiM architectures using power/timing side channels. In contrast, we demonstrate how to establish high-throughput timing covert and side channels using PiM operations.
Other Main Memory-based Timing Attack Defenses. We already extensively discuss defenses that fundamentally eliminate the timing channel (§6) and defenses based on cache management methods that are inapplicable. Another approach is to restrict the fine-grained timers’ usage to prevent attackers from measure memory requests’ latencies, which is applied in some modern processors (e.g., Apple M1 cores [124]). However, these timers are used in many user space applications (e.g., performance and power consumption monitoring, multithreading, and synchronization and locking). Restricting access to these timers can disable these applications or degrade their performance. DAGguise [125] is a defense mechanism against memory timing side-channel attacks that obfuscates an application’s memory access pattern. DAGguise does not mitigate row buffer-based timing channels and requires a closed row policy. Trusted execution environments (TEEs) (e.g., Intel SGX [126, 127]) can isolate applications from the rest of the system. However, prior works already showcase multiple side-channel attacks on TEEs leveraging the main memory [128], and other structures [129, 130, 131, 128, 132, 133, 134, 135, 136, 137].
Enhancing System Security using PiM. Several prior works enhance the security by leveraging PiM. SecNDP [138] proposes an encryption and verification mechanism for untrusted PiM devices. It only focuses on protecting confidentiality and does not consider or analyze potential timing channels. Nelson et al. [139] propose offloading security critical tasks to near-memory compute units to avoid storing sensitive information in caches and eliminate cache-based side-channel attacks. In this work, we demonstrate and evaluate the implications of PiM on main memory-based timing attacks. InvisiMem [140] and ObfusMem [141] propose leveraging 3D-stacked memory to design secure processors that provide oblivious RAM equivalent guarantees.
8 Conclusion
We introduce IMPACT, a set of high-throughput main memory timing attacks that leverage the direct memory access enabled by PiM architectures. IMPACT exploits the shared DRAM row buffer to 1) establish covert channels to communicate with other processes at high throughput and 2) leak security-critical information of victim processes based on their memory accesses. IMPACT covert channels leverage parallelism and eliminate cache bypassing steps, resulting in increased throughput compared to existing main memory-based covert channels. IMPACT side-channel attack targets a genomics application and leaks the private characteristics of a user’s sample genome by leveraging PiM operations. We evaluate and discuss three fundamental defense mechanisms against IMPACT.