### v4

# Revisiting Main Memory-Based Covert and Side Channel Attacks in the Context of Processing-in-Memory

Abstract
We introduce IMPACT, a set of high-throughput main memory-based timing attacks that leverage characteristics of processing-in-memory (PiM) architectures to establish covert and side channels. IMPACT enables high-throughput communication and private information leakage by exploiting the shared DRAM row buffer. To achieve high throughput, IMPACT (i) eliminates expensive cache bypassing steps required by processor-centric memory-based timing attacks and (ii) leverages the intrinsic parallelism of PiM operations. We showcase two applications of IMPACT. First, we build two covert channels that leverage different PiM approaches (i.e., processing-near-memory and processing-using-memory) to establish high-throughput covert communication channels. Our covert channels achieve 8.2 Mb/s and 14.8 Mb/s communication throughput, respectively, which is 3.6$\times$ and 6.5$\times$ higher than the state-of-the-art main memory-based covert channel. Second, we showcase a side-channel attack that leaks private information of concurrently-running victim applications with a low error rate. Our source-code is openly and freely available at https://github.com/CMU-SAFARI/IMPACT.
11footnotetext: F. N. Bostancı and K. Kanellopoulos are co-primary authors.
1 Introduction
Data movement between computation units (e.g., CPUs, GPUs, ASICs) and main memory (e.g., DRAM) is a major performance and energy bottleneck in current processor-centric computing systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22], and is expected to worsen due to the increasing data intensiveness of modern applications, e.g., machine learning [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33] and genomics [34, 35, 36, 37, 38, 39, 40, 41, 42]. To mitigate the overheads caused by data movement, various works propose Processing-in-Memory (PiM) architectures [13, 43, 8, 44, 45, 46, 11, 47, 19, 48, 1, 2]. There are two main approaches to PiM [49, 10, 8]: (i) Processing-near-Memory (PnM) [50, 51, 52, 53, 37, 54, 55, 38, 56, 57, 21, 58, 59, 60, 26], where computation logic is added near the memory arrays (e.g., in a DRAM chip, next to each bank, or at the logic layer of a 3D-stacked memory [61, 62, 63, 64, 65, 66, 20, 67, 21, 11, 68, 25, 69, 35, 26]); and (ii) Processing-using-Memory (PuM) [70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85], where computation is performed by exploiting the analog operational properties of the memory circuitry.
Prior works from industry and academia demonstrate the large performance and energy benefits of PiM architectures. A set of PiM techniques are already implemented in real products [86, 60, 87, 88, 89] and many more are expected to be adopted in the near future. Therefore, it is important to investigate the security aspects of such techniques and implement countermeasures to avoid potential widespread security vulnerabilities that their deployment can cause. Unfortunately, no prior work analyzes and evaluates the security of emerging PiM architectures against timing covert- and side-channel attacks.
In this work, we analyze PiM architectures and show that the adoption of PiM architectures creates opportunities for critical main memory-based timing attacks due to two reasons. First, to eliminate data movement, PiM architectures provide direct access to main memory, which is a key building block for high-throughput main memory-based timing attacks. Second, defenses against these attacks either incur high performance overheads or are not applicable to PiM architectures.
1. Direct Access to Main Memory. Main memory-based timing attacks establish covert communication channels or leak secrets by observing and changing the shared main memory states (e.g., DRAM row buffer states). To achieve high throughput, timing attacks require fast and reliable access to main memory. In today’s systems, it is difficult to access main memory with high frequency due to the deep cache hierarchies that filter the memory requests and incur additional latency overheads [90]. [^1] We provide analysis as to why prior attack primitives that bypass caches do not provide fast and reliable enough direct main memory access in §3.2.
The adoption of PiM techniques changes how applications and computation units access main memory: (i) due to area and thermal dissipation limitations, compute units in PnM architectures have small or no caches that allow the userspace applications to bypass caches more easily (i.e., with fewer instructions and lower latency), and (ii) PuM architectures provide the ability to offload to main memory specialized PiM instructions (e.g., bulk data copy and initialization [75, 91, 73, 92, 85], and bitwise/arithmetic operations [76, 93, 73, 79, 78, 58, 85, 81, 80, 82]) that completely bypass the deep cache hierarchies of modern processors. As a result, PiM architectures provide userspace applications with fast and reliable direct main memory access that can be exploited by attackers for high-throughput timing covert- and side-channel attacks.
2. Hard-to-Mitigate with Practical Defenses. A fundamental solution to memory-based timing attacks is to eliminate the timing channel completely (e.g., by enforcing constant-time memory accesses or introducing intrusive and restrictive memory partitioning between different security contexts). However, eliminating the timing channel incurs very high performance overheads (as shown in Section 7). Therefore, several prior works [94, 95, 96, 97] propose defenses and detection mechanisms that restrict the use of timing channels with cache management-based methods, such as restricting the userspace applications’ access to cache management instructions (e.g., clflush instruction in x86 systems [94]) and detecting timing attacks based on cache access patterns (e.g., a high number of cache misses can indicate a timing attack) [95, 96, 97]. These defenses are inapplicable to PiM architectures because these architectures provide direct access to memory by bypassing the cache hierarchy.
Based on our analysis, we introduce IMPACT, a set of high-throughput In-Memory Processing-based timing Attacks that leverage direct and fast main memory accesses enabled by PiM architectures. Using direct access to DRAM via PiM operations, IMPACT observes and exploits the shared DRAM row buffer states. We demonstrate that IMPACT achieves high communication and information leakage throughput by (i) eliminating expensive cache bypassing steps required by other main memory- and cache-based timing covert- and side-channel attacks, and (ii) leveraging the intrinsic parallelism of PiM operations. We showcase this finding on two covert- and one side-channel attacks. To protect future systems against IMPACT, we describe and analyze four defense mechanisms.
IMPACT-PnM Covert Channel Variant. We build a covert channel exploiting a PnM architecture called PiM-enabled instructions [67] that enables executing instructions in compute units near DRAM. In this variant, the receiver uses PiM-enabled instructions to measure the time it takes to execute a simple arithmetic operation on a set of DRAM rows and detect row buffer conflicts. The sender transmits a message by encoding the message with row buffer conflicts in selected addresses and carefully creating conflicts with the receiver’s accesses using PiM-enabled instructions. By executing instructions near memory, IMPACT-PnM achieves 8.2 Mb/s communication throughput, which is 3.6$\times$ faster than the state-of-the-art main memory-based covert-channel attack [98].
IMPACT-PuM Covert Channel Variant. We build a covert channel exploiting a PuM technique called RowClone [75] that enables bulk data copy and initialization. In this variant, the sender transmits a message with in-memory row-copy operations by copying rows in different DRAM banks in parallel for different bits of the message. The receiver decodes the message by issuing row-copy operations and measuring their latency. By exploiting the parallelism in memory, IMPACT-PuM achieves a communication throughput of 14.8 Mb/s, which is 6.5$\times$ faster than the state-of-the-art [98].
IMPACT Side-Channel Attack. We build a PnM-based IMPACT side-channel attack that exploits PiM-enabled instructions [67] to leak private information of a concurrently-running application. We showcase this side-channel attack on genomic read mapping, one of the fundamental applications in DNA sequence analysis where preserving the privacy of the human genomic data is crucial [34, 99, 36, 100, 101, 102, 103]. Our side-channel attack leverages PiM-enabled instructions [67] to leak the private characteristics of a user’s query genome by observing the memory access patterns of the read mapping application. Our evaluation demonstrates that our side-channel attack leaks the properties of a query genome at a throughput of 7.6 Mb/s with 96% accuracy.
Defense Mechanisms. Based on our security analysis, we discuss and evaluate potential defense mechanisms to mitigate IMPACT attacks by eliminating the timing channel and reducing the attack throughput. We describe the performance versus security trade-offs of these defenses. We eventually conclude that mitigating IMPACT incurs high performance overheads and more research is needed to find low-overhead solutions.
This paper makes the following contributions:
  • To our knowledge, this is the first work that analyzes and evaluates the security of emerging PiM architectures against timing covert-channel and side-channel attacks.
  • We introduce IMPACT, a new set of high-throughput main memory-based timing attacks that leverage PiM operations to gain fast and reliable access to main memory.
  • We demonstrate two IMPACT covert channel variants, IMPACT-PnM and IMPACT-PuM that exploit PnM and PuM techniques, respectively. IMPACT-PnM allows the sender and the receiver to exchange messages using PiM-enabled instructions. IMPACT-PuM further leverages the parallelism provided by bulk in-memory copy operations like RowClone. We show that both attacks outperform the state-of-the-art main-memory-based covert-channel attacks in terms of communication throughput.
  • We demonstrate an IMPACT side-channel attack on genomic read mapping by leveraging PnM, using PiM-enabled instructions. We show that our side-channel attack leaks sensitive information with high throughput and a low error rate.
  • We describe and evaluate the trade-offs of four defense mechanisms to protect PiM-enabled systems against IMPACT’s covert- and side-channel attacks. We call for future research to find low-overhead solutions.
2 Background
This section provides a concise overview of 1) DRAM organization and operation, 2) processing-in-memory, and 3) main memory-based timing attacks.
2.1 DRAM Organization and Operation
Organization. Fig. 1 shows the hierarchical organization of a modern DRAM-based main memory. A memory channel connects the processor (CPU) to a DRAM module, where a module consists of multiple DRAM ranks. A rank consists of multiple DRAM chips that operate in lockstep. Each DRAM chip contains multiple DRAM banks that can be accessed independently. A DRAM bank is organized as a two-dimensional array of DRAM cells, where a row of cells is called a DRAM row. A DRAM cell consists of 1) a storage capacitor, which stores one bit of information in the form of electrical charge, and 2) an access transistor, which connects the capacitor to the row buffer through a bitline controlled by a wordline.
Figure 1: Organization of a typical modern DRAM module.
Operation. To perform a read or write operation, the memory controller first needs to open a row, i.e., copy the data of the cells in the row to the row buffer. To open a row, the memory controller issues an activation command to a bank by specifying the row address to open. After activation is complete, the memory controller issues either read or write commands to read or write a DRAM word within the activated row. Subsequent accesses to the same row can be served quickly without fetching the data to the row buffer, creating a row hit [104]. To access data from another DRAM row in the same bank, the memory controller must first close the currently open row by issuing a bank precharge command, creating a row conflict/miss.
2.2 Processing-in-Memory
Processing-in-Memory (PiM) moves computation closer to where the data resides, alleviating the data movement bottleneck between memory and the processor [13, 43, 8, 44, 45, 46, 11, 47, 19, 48, 1, 2]. There are two main PiM approaches [49, 10, 8]: 1) Processing-near-Memory (PnM) and 2) Processing-using-Memory (PuM). PnM places computation logic near memory arrays (e.g., in a DRAM chip, next to each bank, or at the logic layer of 3D-stacked memory [50, 51, 52, 53, 37, 54, 55, 38, 56, 57, 21, 58, 59, 60, 26]). The PnM logic can execute application regions (e.g., instructions [67, 53], functions [48], application threads [20, 77]) depending on the design. PnM computation units typically employ small caches due to 1) area constraints and 2) the ineffectiveness of caches in improving the performance of a wide range of memory-intensive applications. Therefore, various systems [53, 67, 48] offload the memory-intensive application regions on PnM units, while executing more cache-friendly application regions on the host CPU with a deeper cache hierarchy. PuM uses analog operational properties of the memory circuitry (e.g., DRAM) to enable massively parallel in-memory computation [70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85].
2.3 Main Memory Timing Attacks
Microarchitectural timing attacks exploit the observable microarchitectural state changes to propagate messages between processes and leak secrets. These attacks measure the time to complete an operation that depends on a shared system structure (e.g., accessing a memory location by going through the shared caches) to infer the state of the shared structure.
Main memory timing attacks exploit several shared structures in the main memory. One such structure is the DRAM row buffer that acts as an internal cache inside DRAM banks. An attacker can understand whether a row is recently accessed (i.e., already in the row buffer) by measuring the time it takes to access that row. Using this observation, DRAMA [98] builds a side-channel attack to leak keystrokes in a real system. Several prior works (e.g., [98, 105, 106, 107, 108]) leak DRAM address mapping functions. Other main memory-based attacks exploit the memory bus contention [109] and memory deduplication [110, 111] and (de)compression [112, 113, 114, 115, 116, 117, 118, 119, 120] latencies.
3 A Case for PiM-based Timing Attacks
The adoption of PiM architectures has been gaining momentum in recent years because they provide high performance and low energy consumption by reducing data movement. Therefore, it is timely and critical to analyze and evaluate the security of PiM architectures to prevent any vulnerabilities from causing widespread susceptibility as PiM architectures become widely available. Unfortunately, no prior work analyzes and evaluates the security of emerging PiM architectures against timing covert- and side-channel attacks.
After a thorough analysis of existing PiM architectures, we show that adopting PiM solutions creates opportunities for high-throughput main memory-based timing attacks for two main reasons. First, the adoption of PiM solutions provides userspace applications with fast and reliable direct access to main memory to reduce data movement. Second, defenses against these attacks either incur high performance overheads or are not applicable to PiM architectures. In this work, we study the impact of PiM architectures on row buffer-based timing channels and their throughput.
3.1 Row Buffer Timing Channel
A DRAM row buffer acts as a direct-mapped cache inside a DRAM bank and is shared across applications that can access main memory. Accessing a row that is already in the row buffer results in a significantly lower latency because the memory controller does not need to issue additional commands (i.e., precharge the bank and fetch the target row). Therefore, a userspace application can infer whether a row has been accessed recently by measuring the time it takes to access the row.
To demonstrate the latency difference between a row buffer hit and a row buffer conflict in a PiM-enabled system, we simulate such a system (described in detail in §5.1). We run a microbenchmark that 1) creates memory accesses that result in row buffer hits and conflicts, and 2) measures the memory access latency for each memory request. We observe that a row buffer conflict takes 74 CPU cycles more than a hit for a CPU operating at 2.6 GHz, which is large enough to detect, which aligns with the observations of prior works [98, 107, 108]. Based on this observation, we conclude that a userspace application can exploit the row buffer as a timing channel in a PiM-enabled system.
3.2 Existing Main Memory Attack Primitives
Main memory timing attacks are powerful as they can be used to gather information on all applications that share main memory. However, these attacks have limited throughput in traditional architectures because accessing main memory directly is difficult due to deep cache hierarchies.
There are four main attack primitives that enables attackers to bypass the cache hierarchy and directly access main memory: 1) using specialized instructions to flush a target cache line (e.g., clflush [121]) from the cache deterministically, 2) using cache eviction sets [122] to replace a target cache line in the cache (i.e., by creating many conflicting memory requests to replace the cache line with newly-fetched cache lines), 3) using a (remote) direct memory access engine (i.e., (R)DMA engine) [123, 124], and 4) using non-temporal memory hints (e.g., movnti instruction [121]).
We analyze the attack primitives in terms of efficiency and effectiveness. To do so, we identify four properties of an efficient (i.e., low-latency) and effective attack primitive: 1) avoiding cache lookup overhead, 2) avoiding high-latency memory accesses, 3) creating an easily detectable timing difference, and 4) reliable functionality (e.g., guaranteed to work by the ISA). Table 1 shows a comprehensive comparison of these four attack primitives and PiM operations in terms of these properties.
Table 1: Efficiency and Effectiveness of Attack Primitives
Low Latency Effectiveness
No Cache
Lookup
No Excessive
Memory Accesses
Timing Difference
Detectability
ISA
Guarantees
Specialized
Instructions
Eviction Sets
DMA/RDMA N/A
Non-temporal
Memory Hints
PiM Operations
First, specialized cache flush instructions have two main disadvantages that limit their effectiveness at directly accessing main memory: 1) they put the write-back (to the main memory) latency of a cache line on the critical path to access a memory location [121], and 2) in some modern systems, these instructions are privileged and cannot be used by userspace applications (e.g., ARM processors [125]) (not shown in the table). Second, to directly access main memory, cache eviction sets require many load requests proportionally to the number of ways in the caches, which increases with modern designs. Therefore, the attacks that use them have limited throughput. Even though the instructions used to create the eviction sets are part of the ISA, the eviction set is not guaranteed to evict the target cache line due to the cache replacement policy, cache prefetchers, and other memory system design choices that are not transparent to the attacker. Third, DMA and RDMA engines [123, 124] incur high overheads due to deep software stacks. As a result, attacks using this primitive need to utilize states with larger timing differences to be distinguishable. An example work using this primitive [126] showcases a covert-channel attack based on memory contention. However, memory contention-based timing differences is not as fine-grained as row buffer-based timing differences. Therefore, the communication throughput of these covert channels is limited (in the order of Kb/s). Fourth, the implementation of non-temporal memory hints varies across systems, as system designers are not bound by the instruction set architecture. These instructions may prefetch data from the last level cache (prefetchnta), or can be cached in intermediate buffers (e.g., movnti instruction) [127]. Consequently, the ISA does not guarantee that non-temporal hints will consistently and directly access main memory. In contrast, PiM operations provide fast and reliable direct memory access to userspace applications, which can be exploited for high-throughput covert- and side-channel attacks, as we show in this paper.
3.3 High Throughput Main Memory-based Covert Channels
To demonstrate the impact of the deep cache hierarchy latencies (i.e., cache access and eviction latencies) on an attack’s throughput, we simulate a PiM-enabled system (described in §5.1), establishing a state-of-the-art row buffer-based covert channel [98].
3.3.1 Attack Scenarios
We simulate two attacks. First, we evaluate a baseline main memory attack that relies on cache bypassing using cache eviction sets. To transmit messages through the row buffer, the sender and receiver processes 1) evict the corresponding cache line to directly access main memory and 2) fetch the target row into the row buffer using an additional memory request. The baseline attack evicts a cache line by issuing N memory requests to the target cache set in an N-way cache. The actual eviction latency in a real system can be much higher due to a larger number of memory accesses required to bypass more complex replacement policies employed in today’s systems [128, 129] or due to address translation overheads [130, 131, 132, 133, 134, 135, 136, 137]. Second, we evaluate a direct memory access attack that issues only one memory request to fetch the data to the row buffer without any cache access or evictions.
For both attacks, we measure the time it takes to send a bit through the row buffer, including cache access and eviction latencies, encoding logic-1 values as row buffer hits and logic-0 as row buffer conflicts. We calculate the communication throughput based on the average time to propagate bits from the sender to the receiver. We analyze the impact of (i) the LLC size and (ii) the number of LLC ways on communication throughput to show the limitations in current and future systems with different cache designs.
3.3.2 Impact of LLC Size on Communication Throughput
Fig. 2 shows the impact of the LLC size on (i) the throughput of the baseline and direct memory access attacks on the primary y-axis (left) and (ii) the latency of a cache line eviction on the secondary y-axis (right). The x-axis shows the cache configuration with increasing last-level cache (LLC) sizes and a fixed number of ways (i.e., 16). We calculate the cache access latency with increasing LLC sizes using CACTI 6.0 [138].
Figure 2: Impact of the LLC size on data leakage throughput of covert channels and LLC eviction latency.
We make two key observations. First, the communication throughput of the direct memory access attack is $11.27$ Mb/s across all LLC sizes, which is significantly greater than the throughput of up to $2.29$ Mb/s achieved by the baseline cache eviction sets based attack. Second, the baseline attack’s throughput decreases with increasing LLC sizes since the cache access and eviction become progressively more expensive. With increasing LLC size, the cache access latency increases, amplifying the latency of eviction, which requires multiple cache accesses. Therefore, we conclude that with increasing LLC sizes, cache access and eviction latencies increase, hindering the baseline attack’s communication throughput, compared to the direct memory access attack, which is independent of the LLC size.
3.3.3 Impact of LLC Associativity on Communication Throughput
Fig. 3 shows the impact of LLC associativity on (i) the communication throughput of baseline and direct memory access attacks on the primary y-axis (left) and (ii) the eviction latency of a cache line on the secondary y-axis (right). The x-axis shows the increasing number of LLC ways (from 2 to 128) with a fixed total LLC size (e.g., 16 MB). We observe that as the number of LLC ways increases, the number of memory requests required to evict a cache line increases, greatly reducing the baseline attack’s communication throughput.
Figure 3: Impact of LLC associativity on the data leakage throughput of covert channels and the eviction latency.
Based on these two experiments, we conclude that attacks that communicate with main memory through caches are greatly limited in throughput due to expensive cache access and eviction latencies, which worsen with increasing size and associativity. In contrast, a covert channel based on direct memory access can achieve high throughput independently of the cache configuration, as it does not require bypassing the cache hierarchy.
4 IMPACT: A Set of High-Throughput Main Memory Timing Attacks
We introduce IMPACT, a set of high-throughput main memory-based timing attacks that take advantage of the direct memory access enabled by emerging PiM architectures. IMPACT achieves high communication and information leakage throughput by (i) eliminating expensive cache bypassing steps required by other main memory- and cache-based timing covert- and side-channel attacks and (ii) leveraging the intrinsic parallelism of PiM operations. In this work, we present two covert-channel and one side-channel attacks enabled by PnM and PuM solutions.
4.1 PnM-based IMPACT Covert-Channel Attack
Baseline PnM Architecture. PiM-Enabled Instructions (PEI)  [67] is a PnM architecture that expands the instruction set architecture with instructions that can be executed in compute units in the processor core or near memory based on a dynamic decision made considering the locality of the required memory accesses. We choose PEI as the baseline PnM architecture as it introduces a simple and effective model for PnM, and our attack can be generalized for other PnM architectures with similar design components (e.g., FIMDRAM [89]).
The baseline PEI architecture has two key components. First, it integrates a PEI Computation Unit (PCU) near each DRAM bank and inside the CPU. In-memory PCUs are shared among all cores of the host processor and are responsible for executing PEIs near the respective DRAM bank. Second, there is a PEI Management Unit (PMU) that monitors data locality in application regions with a locality monitor, and decides where to map each region (either to host-side or memory-side PCUs). This way, the PEI architecture can execute application regions with high data locality on the host side to benefit from the cache hierarchy. In contrast, if an application region does not benefit from the cache hierarchy, it is executed on the memory-side PCU to benefit from the lower memory access latency.
Attack Overview. Before the attack, the sender and the receiver co-locate their data in the same set of DRAM banks. To achieve this, one process uses memory massaging techniques to place its data in a page located in the same DRAM bank as the other process, as used in many prior works (e.g., [106, 98, 139, 140]). To incentivize the PnM architecture to execute the instruction in a memory-side PCU and directly access DRAM, the attack processes access different cache lines in the allocated rows, and allocate and use multiple rows in each bank.
IMPACT-PnM uses PEI operations and exploits the latency difference between a DRAM row buffer hit and a conflict in a shared DRAM bank to exchange messages between the sender and the receiver. The sender transmits bits of the message to the receiver in multiple batches of M bits, where the value of each bit is encoded as interference in the row buffer of a DRAM bank. In this attack, we encode logic-1 as interference (i.e., a row buffer conflict), and logic-0 as no interference (i.e., a row buffer hit).
Fig. 4 and Listing 1 show the end-to-end flow of IMPACT-PnM and all the steps involved in establishing the covert channel between the sender and the receiver.
Figure 4: Flow of the PnM-based Covert Channel Attack.
1 Receiver (Step 1):
2 init_row = initialize_DRAM_rows_with_PEIs();
3
4 Sender (Step 2):
5 message[0: N-1]; // N-bit message
6 batch_size = M; // M-bit batch, where M<N
7 for every batch:
8 bank = 0;
9 for each bit in batch_size;
10 if message[bank] == 1: // activate a row
11 pim_add(sender_row[bank]);
12 else: // do not interfere with the receiver
13 NOP();
14 bank++;
15 // ensure all memory requests are served
16 // before starting the next batch
17 memory_fence();
18
19 Receiver (Step 3):
20 for every batch:
21 bank = 0;
22 for each bit in batch_size;
23 timer(start);
24 pim_add(init_row[bank]);
25 timer(end);
26 latency = end - start;
27 if latency > THRESHOLD: // row buffer conflict
28 result = ’1’;
29 else: // row buffer hit
30 result = ’0’;
31 bank++;
32 memory_fence(); // complete the batch
Listing 1: Pseudocode of PnM-based Covert-Channel Attack.
Step 1. The receiver issues PEIs to initialize each DRAM bank by activating a predetermined row (Line 2, 1). Initialization copies the receiver’s data to row buffers.
Step 2. The sender transmits a batch of the message by mapping each bit in the batch to a separate DRAM bank ( 2). Then, the sender checks the bit value (Line 10). If the bit is 1, the sender issues a PEI to activate a row (Line 11, 3) and thus, copies its content to the row buffer ( 4). Otherwise, the sender issues a NOP instruction (Lines 12-13, 5) and does not copy any new data to the row buffer ( 6).
Step 3. The receiver probes the DRAM banks by issuing PEIs to the initialized rows and measuring the PEI latencies (Lines 23-25, 7). If the latency of a PEI exceeds a predetermined row buffer hit threshold, the receiver detects interference and decodes the bit as a 1 (Lines 27-28, 8). Otherwise, the access is considered a row buffer hit, and the receiver decodes the bit as a 0 (Lines 29-30, 9). After the transmission of a batch of the message (i.e., M bits), the sender and the receiver execute a memory fence (Lines 17 and 32) to ensure that all the PEIs are executed before moving on to the next batch.
The receiver accesses the next cache line in the initialized row to ensure its PEI is executed near memory by bypassing the locality monitor. If the number of batches is greater than the number of cache lines in a row, it uses another row in the same bank to measure access latency. The sender and the receiver repeat the same steps for each batch until the entire message is transmitted.
Sender-Receiver Synchronization. The sender and the receiver synchronize to 1) ensure their actions do not interfere with each other and 2) overlap the latencies of their operations to increase the throughput of the attack. In our proof-of-concept implementation, the sender and the receiver use a semaphore. The semaphore’s value indicates the number of batches that the sender has completed transmitting, but the receiver has not yet probed. The receiver blocks on the semaphore until the sender increments it (i.e., does not start probing). When the sender completes the transmission of a batch, they increment the semaphore to signal the receiver to start. Upon receiving the signal, the receiver decrements the semaphore and starts probing the DRAM banks. When the receiver completes probing, they read the semaphore again, and the process repeats until the entire message is transmitted.
4.2 PuM-based IMPACT Covert-Channel Attack
Baseline PuM Architecture. We assume a PuM architecture that provides user applications with RowClone [75] which enables bulk data copy and initialization operations within DRAM’s subarray boundaries by exploiting the internal operations of DRAM. RowClone’s Fast Parallel Mode works by issuing two activation commands back-to-back: the first activation copies the content of a source row (i.e., src) into the row buffer. The second connects the destination row (i.e., dst) to the bitlines. Since the row buffer stores the src’s data after the first activation, by the time the dst is activated, the data in dst is overwritten by src’s data.
To execute RowClone, we assume that the userspace application specifies 1) a source virtual address range, 2) a destination virtual address range, and 3) a mask to copy specific parts of the source range to the destination range. This way, the memory controller receives one RowClone request and breaks it into parallel RowClone requests, one for each set bit of the mask.
Attack Overview. Our PuM-based covert channel attack (IMPACT-PuM) exploits in-DRAM bulk copy operations to establish a covert channel between a sender and a receiver. Before the attack, the sender and the receiver co-locate their data in the same set of DRAM banks. The sender transmits a message to the receiver in multiple batches of M bits. To do so, the sender issues RowClone operations by carefully selecting 1) source and destination address ranges corresponding to physical pages spanning the selected set of DRAM banks, and 2) a mask to cause interference only in selected DRAM banks. The sender and receiver processes synchronize with barriers similar to IMPACT-PnM (§4.1)
Fig. 5 and Listing 2 show the end-to-end flow of IMPACT-PuM and the steps for establishing the covert channel between the sender and the receiver.
Figure 5: PuM covert-channel attack flow.
Step 1. First, the receiver initializes the DRAM banks by issuing a RowClone operation (Line 25, 1).
Step 2. The sender transmits the message by dividing the message into M-bit batches. Then, the sender maps each bit of a batch to a separate DRAM bank and creates a mask for the batch (Line 18), determining which banks will be executing a RowClone operation (Line 18). By using a mask, the sender uses only a single RowClone operation to transmit an M-bit batch of the message (where M is the number of banks) (Line 20, 2). The memory controller issues individual RowClone operations to each DRAM bank based on the specified mask (Lines 1-9). To send a 1, the sender triggers RowClone in the corresponding DRAM bank and activates a different row from the receiver’s row in the bank (Lines 4-7, 3). To send a 0, the sender does not trigger RowClone in the corresponding DRAM bank and does not interfere with the receiver’s activated row (Lines 8-9, 4).
Step 3. The receiver detects interference in DRAM banks using RowClone operations as follows. The receiver issues one RowClone operation for each DRAM bank to precisely measure the time it takes to complete the operation in each bank (Lines 29-33). To detect if the sender activated another row, the receiver uses its initial destination address range as the source (i.e., swapping the direction of the copy operation) (Line 31). If the sender induces interference in a DRAM bank, the receiver observes high latency for the RowClone operation, indicating the corresponding bit in the message is set to 1 (Lines 34-35, 5). Otherwise, the receiver detects a fast RowClone operation, which indicates that the bit is set to 0 (Lines 36-37, 6).
1 Memory Controller:
2 row_clone(src, dst, mask):
3 for each DRAM bank:
4 if mask[bank] == 1:
5 // copy the corresponding src row
6 // to the destination row
7 row_clone(src, dst, bank);
8 else:
9 NOP(); // do not interfere with the receiver
10
11 Sender:
12 message[0:N-1]; // N-bit message
13 batch_size = M; // M-bit batch, where M<N
14 start = 0;
15 for every batch:
16 end = start + batch_size;
17 // create mask for the batch
18 mask = message[start:end];
19 // issue a RowClone operation using the mask
20 row_clone(src, dst, mask);
21 start = end; // update start for the next batch
22 memory_fence(); // complete the batch
23
24 Receiver:
25 init_DRAM_rows_with_RowClone();
26 for every batch:
27 for each DRAM bank:
28 mask = 1 << bank;
29 timer(start);
30 // issue RowClone to one bank at a time
31 row_clone(dst, src, mask);
32 timer(end);
33 latency = end - start;
34 if latency > THRESHOLD: // row buffer conflict
35 result = ’1’;
36 else: // row buffer hit
37 result = ’0’;
38 memory_fence(); // complete the batch
Listing 2: Pseudocode of PuM-based Covert-Channel Attack.
Advantage over IMPACT-PnM (§4.1). In IMPACT-PuM, the sender can use a single RowClone operation to transmit an M-bit message to the receiver in parallel, where M is the number of DRAM banks. Thus, IMPACT-PuM requires fewer computational resources to transmit the same amount of information compared to IMPACT-PnM. For example, using a single thread, the sender can activate all M DRAM banks and transmit M bits of information in parallel. In contrast, in IMPACT-PnM (§4.1), the sender needs to execute one PEI for every bit.
4.3 PnM-based Side-Channel Attack on Genomic Read Mapping
We build a PnM-based side-channel attack that exploits PiM operations to leak private information of a concurrently running application. To demonstrate the IMPACT side-channel attack, we target a DNA sequence analysis application.
Genomic Read Mapping. DNA is a unique identifier of an individual [103], such that even a small fraction of DNA can include sufficient information for linking a DNA sample to the owner’s identity [100, 101]. Therefore, preserving privacy while processing genomic data carries utmost importance to protect sensitive information. Sequencing the DNA of a biological sample produces short DNA fragments called reads [141]. The reads are then processed by read mapping [34, 99, 36, 100, 101, 102, 103], which is a fundamental task performed to find the matching locations of reads in the reference genome that acts as a template to construct the complete genome of the sample.
Read mapping (RM) has two key steps: (i) seeding [142] and (ii) alignment [143, 144]. Fig. 6 shows a high-level overview of RM. During the seeding step, RM identifies the possible locations of the reads in the reference genome based on sequence similarities between the reads and the reference genome. Seeding involves hashing small segments of the DNA sequences (i.e., seeds) ( 1) and probing a hash table that is built from the reference genome ( 2). The hash table provides potential locations for the seed in the reference genome ( 3). In the alignment step, RM aligns the reads with their candidate regions in the reference (i.e., potential reference segments) using dynamic programming methods [99, 100] to find each read’s best matching position among all possible locations ( 4).
Figure 6: DNA Read Mapping Operation: Seeding and Sequence Alignment.
RM requires a large amount of data movement to access different parts of the hash table and corresponding regions on the reference genome. Hence, several prior works aim to improve the performance of RM by offloading seeding and alignment to PiM-enabled systems to reduce the data movement overheads [37, 145, 146, 40, 147, 148, 149, 150, 35, 151, 152]. In this work, we demonstrate that an attacker can launch a high-throughput side-channel attack against RM by utilizing PiM instructions and quickly leak a large amount of information about the query genome.
Attack Overview. We assume that the victim is running an RM application on a PiM-enabled system and the attacker is able to run a malicious application on the same system. The attacker and the victim are using the same read mapping tool (e.g., minimap2 [153]). The read mapping tool constructs a hash table that contains information about the seed locations in the reference genome, and allows each user in the system to probe the hash table via queries during read mapping. We assume the hash table is distributed across multiple DRAM banks. This assumption is realistic, as many modern DRAM address mapping schemes (e.g., [154, 155, 156, 157]) aim to interleave consequent memory chunks across different DRAM banks to exploit bank-level parallelism [158, 159].
Fig. 7 shows the high-level overview of the attack. The hash table is distributed across banks as explained above ( 1). During the attack, the victim application extracts seeds from the reads and offloads the corresponding seeding and alignment steps to the PiM-enabled system ( 2). During the seeding step, the corresponding DRAM row gets activated (i.e., Index 0-M of the hash table) ( 3). The attacker continuously uses PnM instructions to probe the shared hash table in an attempt to identify if the victim application is accessing a specific index of the hash table ( 4). If the attacker observes that the victim application is accessing a specific index, it retrieves partial or exact information about the potential locations of the reads in the reference genome. As shown in previous works on DNA imputation [160, 161, 162, 163], the attacker can use the leaked information in a completion attack [101] to infer properties about some regions of the private query genome ( 5).
Figure 7: High-level overview of the IMPACT attack model on PiM-enabled genomic read mapping.
5 Methodology
5.1 Threat Models
Covert Channel Threat Model. We assume a scenario where a sender and a receiver operate on the same PiM-enabled system to exchange information. PiM instructions are available to userspace applications without any restrictions. The receiver has access to cpuid and rdtscp instructions [121], enabling high-precision measurement of memory access latencies. For IMPACT-PnM, the sender and the receiver co-locate data in the same set of DRAM banks. For IMPACT-PuM, the sender allocates two virtual address ranges whose corresponding physical pages span across the same set of DRAM banks.
Side Channel Threat Model. We assume a scenario where an attacker process operates on the same PiM-enabled system as a victim process and leaks information about the victim process. We use a genomic read mapping implementation based on minimap2 [153], and we modify the implementation to offload the seeding and alignment steps to the PiM-enabled system [67]. We assume that the alignment step includes chaining [99, 153]. We use the human genome as the reference and compare it against synthetic query genomes. We experiment with multiple seed sizes and report the best-performing selection.
5.2 Evaluation Methodology
5.2.1 Proof-of-Concept
We construct a proof-of-concept of all IMPACT variants using the Sniper Multicore Simulator [164]. [^2] There is no commercially available system that follows a similar architecture as PEI or any modern system that enables end-to-end RowClone operations. We extend Sniper to (i) accurately model all internal DRAM operations, (ii) emulate and simulate the PiM interface and PiM operations, and (iii) emulate the functionality of the cpuid and the rdtscp instructions [121] to measure execution time.
Table 2 shows the configuration of our simulation setup. Before launching the attack, the sender and the receiver threads warm up to avoid compulsory DRAM accesses for instructions or page table walks. For the PnM system [67], we model the additional latency of a PEI (e.g., latency of accessing additional system structures) as $3$ cycles [67].
Table 2: Simulated System Configuration
CPU 4-core, 4-way issue, OoO x86 core, 2.6GHz
MMU L1 DTLB (4KB): 64-entry, 4-way, 1-cycle, L1 DTLB (2MB): 32-entry, 4-way, 1-cycle, L2 TLB: 1536-entry, 12-way, 12-cycle
L1 Cache I-cache: 32 KB; L1D Cache: 32 KB, 8-way, 4-cycle, LRU, IP-stride prefetcher [165]
L2 Cache 2 MB, 16-way, 16-cycle, SRRIP [166], Streamer prefetcher [167]
L3 Cache 2 MB/core, 16-way, 50-cycle, SRRIP [166]
Main Memory DDR4-2400, 16 Banks, 4 bank groups, 1 Rank, 1 Channel, Row size = 8192 bytes, $t_{RCD}$ = 13.5 ns, $t_{RP}$ = 13.5 ns , $t_{RC}$ = 13.5 ns, Open Row policy, Row Timeout = 100 ns
5.2.2 Comparison Points for Covert-Channel Attacks
To showcase the effectiveness of IMPACT, we evaluate the leakage throughput of five different covert-channel attacks.
(i) DRAMA-clflush: The row buffer-based covert-channel attack proposed in [98] which uses clflush instructions [121] to send memory requests directly to main memory. clflush only probes the LLC to flush the cache line.
(ii) DRAMA-Eviction: The row buffer-based covert channel attack proposed in [98], which uses cache eviction sets [122] to send memory requests directly to main memory.
(iii) DMA Engine: A row buffer-based attack that uses the DMA engine [123, 124] to bypass the cache hierarchy.
(iv) IMPACT-PnM (§4.1): An attack that bypasses the locality monitor and offloads PEI instructions directly to compute units near DRAM banks in order to establish a covert channel.
(v) IMPACT-PuM (§4.2): An attack that leverages RowClone to offload bulk-copy operations directly to main memory and establish a covert channel.
5.2.3 Noise Sources in Simulation Setup and Error Rate
We simulate hardware prefetchers [167, 165] and page table walkers [121] to induce noise in the simulated system. We measure the throughput of each attack only based on the successfully leaked data (i.e., incorporating error rate in our measurements).
6 Evaluation
6.1 Validation of Proof-of-Concept
We first demonstrate proof-of-concept attacks for IMPACT-PnM and IMPACT-PuM. We send 16-bit messages with a batch size of 16 (i.e., using 16 banks in parallel). Fig. 8 shows the latency (in cycles) measured by the receiver during transmission of example messages. The y-axis shows the measured latency for each (a) PEI and (b) RowClone operation for IMPACT-PnM and IMPACT-PuM, respectively. The x-axis shows the transmitted bit values where each bit is transmitted using a different bank. We make two key observations. First, for both attacks, the receiver successfully determines the transmitted bit values by detecting the row buffer conflicts caused by the sender. Second, the receiver decodes the complete message using a threshold value of 150 cycles. If the measured latency is less than 150 cycles (i.e., row buffer hit), the receiver decodes a bit as logic-0; otherwise (i.e., row buffer conflict), it decodes the bit as logic-1.
Figure 8: Latency (in cycles) measured by the receiver to execute (a) a PEI and (b) a RowClone operation for each bank to decode a 16-bit message.
6.2 Covert Channel Throughput Analysis
Fig. 9 shows the throughput achieved by IMPACT-PnM, IMPACT-PuM and state-of-the-art main memory-based covert-channel attacks in modern systems with increasing LLC sizes. The x-axis shows the increasing LLC size (from 1 MB to 128 MB), and the y-axis shows the leakage throughput of the covert channels. For all attacks, one sender and one receiver process run concurrently and transmit a fixed-length message.
We make four key observations. First, IMPACT-PnM and IMPACT-PuM achieve significantly higher throughput than all other attack vectors. IMPACT-PnM and IMPACT-PuM provide communication throughputs of 8.2 Mb/s and 14.8 Mb/s, respectively, irrespective of the cache size. They support up to 3.6$\times$ and 6.5$\times$ of the throughput achieved by the state-of-the-art main memory-based attack, DRAMA-clflush. Second, IMPACT-PuM provides 82% higher throughput than IMPACT-PnM. This is because the sender in IMPACT-PuM can transmit a message with lower latency using fewer operations (§4.2). Third, as the LLC size increases, the throughputs of DRAMA-eviction and DRAMA-clflush decrease significantly due to the increasing LLC access latency. Fourth, the DMA engine attack provides a throughput of 0.81 Mb/s irrespective of the LLC size, which is 10.1$\times$ slower than IMPACT-PnM due to its OS-related overheads (e.g., context switch overhead, OS instructions). We conclude that IMPACT-PnM and IMPACT-PuM achieve significantly higher throughput than the existing state-of-the-art main memory-based covert-channel attacks.
Figure 9: Leakage throughput of IMPACT-PnM, IMPACT-PuM, and state-of-the-art main memory-based covert channel attacks.
Breakdown of Throughput Improvements. We analyze the throughput breakdown of IMPACT-PnM and IMPACT-PuM by calculating the time spent on sender and receiver routines to transmit a fixed-length message. Fig. 10 shows the time (in CPU cycles) it takes for the sender to send the message and for the receiver to decode it in IMPACT-PnM and IMPACT-PuM. We make two key observations. First, in IMPACT-PuM, the sender routine takes 11.1$\times$ less time than the sender in IMPACT-PnM. This is because IMPACT-PuM can issue only one RowClone request to transmit N bits, where N is the number of banks. In contrast, IMPACT-PnM transmits the same message with potentially multiple PEIs. Second, the receiver routines of IMPACT-PnM and IMPACT-PuM spend similar amounts of time decoding the message as they both need to use distinct PiM operations to detect which DRAM banks were accessed by the sender. Based on these observations, we conclude that IMPACT-PuM has a higher throughput than IMPACT-PnM due to the differences in the sender routines.
Figure 10: Time (in cycles) it takes (i) for the sender to send the message and (ii) for the receiver to read the message in IMPACT-PnM and IMPACT-PuM.
6.3 Side-Channel Attack Throughput Analysis
We evaluate the side-channel attack against genomic read mapping (RM) (§4.3) in a PiM-enabled system [67], sweeping the number of DRAM banks that store the hash table used in the seeding step. Fig. 11 shows the throughput and the error rate of the side-channel attack in the primary y-axis and secondary y-axis, respectively. The x-axis shows the total number of DRAM banks in the system. We measure the throughput of the attack based on the correct guesses of the hash table entries accessed and calculate the error rate based on the number of incorrect guesses. [^3] The exact end-to-end accuracy of identifying the properties of the query genome depends on the chaining mechanism and the imputation algorithm [160, 161] used by the attacker, and it is out of the scope of this work.
Figure 11: Leakage throughput and error rate of the genomic read mapping side-channel attack.
We make two key observations. First, the attacker can leak information about the query genome at a throughput of 7.57 Mb/s and an error rate of $<5$% in a system with 1024 banks. Second, as the number of banks increases, the throughput of the attack decreases to 2.56 Mb/s and the error rate increases to $<15$%. This is because the attacker has to probe more banks, making the attack more prone to noise and reducing its bandwidth.
However, at the same time, the leaked information provides more exact information about the query genome since the number of hash table entries per bank decreases, reducing the number of potential locations of the reads in the reference genome per bank. For example, in a PiM-enabled DRAM device with 1024 banks, the victim activates a row with 16 hash table entries. Assuming the attacker identifies the activated bank correctly, they determine that one of the 16 hash table entries was targeted by the victim. In contrast, in a PiM-enabled device with 2048 banks, the victim activates a row with 8 hash table entries. When the attacker identifies the activated bank correctly, they determine that one of the 8 hash table entries was targeted by the victim, which leaks more precise information on the query genome.
Based on these observations, we conclude that an attacker can successfully leak query genome characteristics in a PiM-enabled system by leveraging PiM operations.
7 Defenses Against IMPACT
We demonstrate high-throughput attack vectors that can emerge with the adoption of PiM architectures. Our results provide insights into new side and covert channel attack vectors that PiM systems might face, and show that these attack vectors should be considered while designing such PiM systems. In this section, we discuss the effectiveness of four defense mechanisms that mitigate the row buffer-based timing channel we introduce.
7.1 Memory Partitioning (MPR)
Bank-level memory partitioning (similarly to [168, 169]) allocates each DRAM bank to only one process and enforces that only that process accesses the DRAM bank. This way, none of the processes can observe latency differences due to other processes’ memory accesses. Thus, the attacker processes cannot build communication channels or leak information. However, this approach has three major drawbacks. First, it limits the number of applications that can run concurrently due to the fixed number of DRAM banks available in the system. Second, it can cause memory underutilization due to allocating DRAM bank-sized memory to applications. Third, it disables data sharing across processes, which can significantly increase data duplication and memory underutilization, and lead to frequent swapping from the disk when the memory space is full.
7.2 Closed Row Policy (CRP)
Enforcing a memory controller policy that closes a DRAM row after each operation, a closed row policy (similar to as described in [155, 170, 171]), can be an effective defense against IMPACT. By closing the DRAM row after each operation, the memory controller ensures that subsequent memory accesses cannot access the same row without requiring a row activation. The closed row policy disrupts the timing patterns exploited by IMPACT. However, it also leads to significant performance overheads [172, 170], even for workloads with low data locality, as every DRAM access leads to a row buffer miss.
7.3 Constant-Time DRAM Accesses (CTD)
Another defense mechanism is to enforce a constant-time DRAM access (CTD) policy. When a memory access is issued, the memory controller waits for a fixed amount of time before returning the data to the processor. This fixed amount of time is equal to the worst-case DRAM access latency. By enforcing CTD, the memory controller ensures that all memory accesses take the same amount of time, which eliminates the timing differences exploited by IMPACT. However, this leads to significant performance overheads (§7.4) as every DRAM access has the worst-case latency.
7.4 Reducing the Attack Throughput with Adaptive Constant-Time DRAM (ACT)
Another approach is to reduce the leakage throughput (without completely getting rid of the covert-channel attack) while incuring lower performance overheads than existing defenses. To understand the effectiveness of such a solution, we implement a memory controller-based countermeasure that adaptively enforces constant-time memory access latency per DRAM bank based on the row buffer contention observed. We call this countermeasure Adaptive Constant-Time (ACT). ACT counts the row buffer conflicts for each DRAM bank during an epoch and decides which latency policy (i.e., constant-time or default latency) to employ in the next one. It switches to the constant-time policy for a bank when there are more row buffer conflicts than a predetermined threshold in the last epoch. This way, in the next epoch, all memory requests being served by the bank will exhibit the same latency (i.e., the worst case latency), and the attacker always observes interference. While employing constant-time latency, ACT still keeps track of the row buffer conflicts. It switches back to the default policy only when the number of conflicts is lower than the threshold in the last epoch. This can reduce the performance overheads caused by CTD. ACT does not fundamentally solve the row buffer-based timing channel, and an attacker can still mount the attack by transmitting messages only when ACT is employing default latency and then waiting while ACT employs constant-time latency. However, this limits the usable epochs during execution and reduces the attack throughput.
We implement ACT and evaluate its impact on system performance and the attack throughput. We evaluate three different configurations an epoch is set to 1000ns: (i) ACT-Aggressive: employing constant-time latency for the next 4000 epochs after the 1st conflict occurs in the bank, (ii) ACT-Mild: employing constant-time latency for the next 2 epochs after the 1st conflict occurs in the bank, (iii) ACT-Conservative: employing constant-time latency for the next 2 epochs after 5 conflicts occur in the bank. To analyze ACT’s impact on system performance, we evaluate four workloads from the GraphBIG benchmark suite [173]: Betweenness Centrality, Breadth-First search (BFS), Connected components (CC), Triangle counting (TC) and, XSBench (XS) [174]. To analyze ACT’s impact on attack throughput, we calculate the reduction in IMPACT-PnM’s throughput. Figure 12 shows the performance comparison of ACT-Conservative, ACT-Mild, ACT-Aggressive, and CTD normalized to a system with no IMPACT defense.
Figure 12: Performance comparison of Constant-Time policy and Adaptive Constant-Time defense policies over the baseline system with no IMPACT defense.
We make two key observations. First, ACT-Aggressive causes similar performance overhead as CTD, while reducing the throughput of IMPACT-PnM by 72% on average (not shown in the figure), in contrast to CTD, which prevents the attack completely. Second, the two less aggressive variants of ACT (ACT-Mild and ACT-Conservative) lead to only 10% performance overhead. However, they cannot reduce the throughput of IMPACT-PnM. This is because the attack utilizes all available banks during a batch of transmission, which increases the idle time between each transmission that uses a specific DRAM bank (i.e., the attack uses the same bank only after a time period where it issues requests to the other banks). This requires significantly high delay times for the mitigation to reduce the throughput. Our findings highlight the necessity for a new approach in defending against main-memory timing attacks.
8 Discussion
8.1 Other Potential PiM-Based Attack Vectors
In this work, we show that emerging PiM architectures can enable a set of high-throughput main memory-based timing attacks based on the DRAM row buffer. Adopting PiM architectures can lead to other attack vectors or exacerbate existing ones, requiring further exploration and analysis. One such example is the impact of PiM on read disturbance vulnerability [175, 172]. A prior PiM proposal in 2021 [82] states that the proposal may lead to increased vulnerability to RowHammer [175] attacks. A new work in 2025, PuDHammer [176], studies the read disturbance effects of Processing-using-DRAM (PuD) operations and shows that these operations greatly increase read disturbance vulnerabilities of DRAM modules (up to 158.6$\times$). Thus, it is critical to employ robust and scalable RowHammer [175] and RowPress [172] solutions to prevent read disturbance bitflips in PiM-enabled systems.
Other potential attack vectors include 1) thermal attacks that leverage PiM operations to increase the temperature in DRAM to make DRAM unavailable and restrict its service time, and degrade reliability, and 2) timing attacks that leverage coarse-grained function offloading. Exploring these directions requires special attention and research, which we leave for future work.
8.2 Applicability of IMPACT and its Defenses
to Complex PiM Architectures
Our observations apply to other PiM architectures as long as they satisfy the key properties that enable the attack: 1) allowing co-locating data from different applications in a DRAM bank, and 2) enabling subsequent accesses from different applications targeting the same DRAM bank. Complex PnM architectures (e.g., systems like Tesseract [177, 178, 179] that employ sophisticated prefetching and network functionality) can degrade the attack throughput by introducing additional noise.
The performance overhead of an IMPACT defense heavily depends on how the defense mechanism interacts with the underlying system. The defense mechanisms we discuss in §7 are not specifically tailored for a given underlying system (e.g., constant-time policy). The new defense mechanism that we introduce in §7.4 enables the constant-time policy only for the bank under contention, reducing the performance overhead, but does not provide strong security guarantees.
8.3 Restricting Access to PIM Operations
A key contribution of our study is to show that unrestricted access to PiM operations enables potentially harmful high-throughput side and covert channel attack vectors. None of the current PiM proposals we are aware of discuss restricting access to PiM operations to ensure security. We hope and believe that our study will guide system designers in balancing the performance, flexibility, and security trade-offs while implementing effective access control mechanisms to PiM.
8.4 Applicability to Future DRAM Devices
The row buffer timing channel would persist in future DRAM-based systems, enabling IMPACT. With newer generations, we observe that DRAM modules consist of increasingly more banks, which increases IMPACT’s covert channel throughput as the attack would benefit from more parallelism. Another change is the introduction of RowHammer mitigations, PRAC [180, 181, 182] and RFM [183], in recent DDR specifications that block DRAM accesses when necessary to prevent RowHammer bitflips. These mitigations incur at least 350ns (at most 1400ns) latency per preventive action. This is significantly higher than the row buffer conflict latency and can be filtered out by the receiver.
We believe that increasing DRAM’s complexity will increase attack vectors as manufacturers try to combat the challenges of technology scaling. For example, recent preprints show that PRAC and RFM introduce timing channels that can be exploited for side-channel [184] and covert-channel [184, 185] attacks.
9 Related Work
To our knowledge, this is the first work that evaluates timing attack vulnerabilities in PiM systems.
9.1 Security Implications of PiM Systems
A set of prior works [186, 187, 188] discuss specific security vulnerabilities in PiM systems. Arafin et al. [186] survey and theoretically analyze security vulnerabilities, focusing on five specific attacks. In this work, we quantitatively and qualitatively analyze and evaluate main memory-based covert and side channels in detail. Wang et al. [188] extract neural network architectural information from an RRAM accelerator using power side channels. Ensan et al. [187] propose an attack on RRAM-based PiM architectures using power/timing side channels. In contrast, we demonstrate how to establish high-throughput timing covert and side channels using PiM operations.
9.2 Other Timing Attacks
We already discuss main memory-based timing attacks in §2.3. In addition, prior works showcase multiple cache-based timing side and covert-channel attacks (e.g., [189, 190, 191, 192, 193, 194, 195, 196, 197]). Among these works, Streamline [192] has the highest throughput as it avoids explicit cache flushes and utilizes a large transmission set to evict cache lines implicitly by thrashing the cache. IMPACT achieves a leakage throughput within the same order of magnitude of Streamline and does not require a transmission set larger than the cache size.
9.3 Other Main Memory Timing Attack Defenses
We already extensively discuss defenses that fundamentally eliminate the timing channel (§7). Another approach is to restrict the high-resolution timers’ usage to prevent attackers from measuring memory requests’ latencies, which is applied in some modern processors (e.g., Apple M1 cores [198]). However, these timers are used in many userspace applications (e.g., performance and power consumption monitoring, multithreading, and synchronization and locking). Restricting access to these timers can disable these applications or degrade their performance. DAGguise [199] is a defense mechanism against memory timing side-channel attacks that obfuscates an application’s memory access pattern. DAGguise does not mitigate row buffer-based timing channels and requires a closed row policy. Trusted execution environments (TEEs) (e.g., Intel SGX [200, 201]) can isolate applications from the rest of the system. Prior works [202, 203, 204, 205, 206, 207, 208, 209, 210, 211] already demonstrate multiple side-channel attacks on TEEs leveraging main memory [205], cache hierarchy [202, 203, 204, 205, 206], page table entries [207, 208, 209, 205], and branch predictors [210, 211].
9.4 Enhancing System Security using PiM
Several prior works enhance system security by leveraging PiM. SecNDP [212] proposes an encryption and verification mechanism for untrusted PiM devices. It only focuses on protecting confidentiality and does not consider or analyze potential timing channels. Nelson et al. [213] propose offloading security-critical tasks to PnM units to avoid storing sensitive information in caches and eliminate cache-based side-channel attacks. In this work, we demonstrate and evaluate the implications of PiM on main memory-based timing attacks. InvisiMem [214] and ObfusMem [215] propose leveraging 3D-stacked memory to design secure processors that provide oblivious RAM equivalent guarantees.
10 Conclusion
We introduce IMPACT, a set of high-throughput main memory timing attacks that leverage the direct memory access enabled by Processing-in-Memory (PiM) architectures. IMPACT exploits the shared DRAM row buffer to 1) establish covert channels to communicate with other processes at high throughput and 2) leak security-critical information of victim processes based on their memory accesses. IMPACT covert channels leverage parallelism and eliminate cache bypassing steps, resulting in increased throughput compared to existing main memory-based covert channels. IMPACT side-channel attack targets a genomics application and leaks the private information with a low error rate. We evaluate and discuss four potential defense mechanisms against IMPACT and eventually conclude that mitigating IMPACT incurs high performance overhead and more research is needed to find low-overhead solutions.
Acknowledgments
We thank the anonymous reviewers of USENIX Security’24, ISCA’24, MICRO’24, ASPLOS’25, and DSN’25 for their constructive feedback. We thank the SAFARI Research Group members for providing a stimulating intellectual environment. We acknowledge the generous gifts from our industrial partners, including Google, Huawei, Intel, and Microsoft. This work, along with our broader work in Processing-in-Memory and memory systems, is supported in part by the Semiconductor Research Corporation (SRC), the ETH Future Computing Laboratory (EFCL), AI Chip Center for Emerging Smart Systems (ACCESS), sponsored by InnoHK funding, Hong Kong SAR, European Union’s Horizon programme for research and innovation [101047160 - BioPIM], a Google Security and Privacy Research Award, and the Microsoft Swiss Joint Research Center.
Appendix A Artifact Appendix
A.1 Abstract
This artifact provides the source code and scripts to reproduce the experiments and results presented in our DSN 2025 paper. It includes implementations of various covert channel attacks (PNM, DRAMA-clflush, DRAMA-evict, PUM, DMA) and a defense mechanism, along with scripts to automate the setup, execution, and result analysis.
A.2 Artifact Checklist (Meta-information)
  • Program: C++ programs, Python scripts, shell scripts.
  • Compilation: GNU Make.
  • Run-time environment: Linux (tested on Ubuntu 20.04 and 22.04), Python 3.
  • Execution: Bash scripts, Python scripts.
  • Metrics: Timing results, performance metrics of attacks and defenses.
  • Output: Figures in PDF format and related data in plaintext files.
  • How much disk space required (approximately)?: 20GB.
  • How much time is needed to prepare workflow (approximately)?: $\sim 1$ hour.
  • How much time is needed to complete experiments (approximately)?: 2-4 hours.
  • Publicly available?: Yes
  • Archived (provide DOI)?: Yes, DOI: 10.5281/zenodo.15116851
A.3 Description
A.3.1 How to Access
The source code and scripts can be downloaded from Zenodo (https://zenodo.org/records/15116851) and GitHub (https://github.com/CMU-SAFARI/IMPACT).
A.3.2 Hardware dependencies
  • x86-64 system.
  • At least 8GB of RAM (128GB recommended for parallel defense experiments).
  • $>64$ cores are suggested for parallel defense experiments.
A.3.3 Software dependencies
  • Linux Operating System (tested on Ubuntu 20.04 and 22.04).
  • Bash.
  • Python (version 3.12.9)
  • GNU Make (version 4.3).
  • wget, tar.
  • GCC (11.4.0), g++ (11.4.0).
  • zlib1g-dev.
A.3.4 Data sets
The defense mechanism experiments may require trace files. The script includes commented-out lines to download example traces. You may need to download these or provide your own traces.
A.4 Installation
A.4.1 Dependency Installation
To install all dependencies, run the dependency installation script: install_dependencies.sh.
{mdframed}
[backgroundcolor=darkgray, linecolor=black, linewidth=0.5pt] \$ cd IMPACT
\$ sh install_dependencies.sh
\$ bash
\$ conda activate impact
A.4.2 Conda Environment Setup
If you would like to install Miniconda manually, follow the steps below.
  1. 1.
  2. 2.
    Run the Miniconda installer.
  3. 3.
    Open a new terminal.
A.5 Experiment Workflow
To run all experiments, run the main script: artifact.sh. This script 1) builds the attacks and the simulator and 2) runs all experiments required to reproduce the key results.
{mdframed}
[backgroundcolor=darkgray, linecolor=black, linewidth=0.5pt] \$ sh artifact.sh
The artifact.sh script will execute the following steps.
A.5.1 Building the Attacks and Simulator
  1. 1.
    The script will navigate to the impact/covert_channel_attack/ directory.
  2. 2.
    It will clean and build the attack code using make clean and make all.
  3. 3.
    It will navigate to the simulator/sniper directory.
  4. 4.
    It will clean and build the simulator using make distclean and make -j\$(nproc).
A.5.2 Running Examples and Attacks
A.5.3 Basic Example Run
The script includes a commented-out line to run a basic example using the Sniper simulator. Uncomment the line.
A.5.4 Running Attacks for Figure 9
The script executes the following attacks:
  • PNM: python3 \$home_dir/impact/covert_channel_attack/script/run_pnm.py \$home_dir 16 \$home_dir/results_attacks/pnm
  • DRAMA (clflush): python3 \$home_dir/impact/covert_channel_attack/script/run_drama.py \$home_dir 16 \$home_dir/results_attacks/drama
  • DRAMA (eviction sets): python3 \$home_dir/impact/covert_channel_attack/script/run_drama_ev.py \$home_dir 16 \$home_dir/results_attacks/drama_ev
  • PUM: python3 \$home_dir/impact/covert_channel_attack/script/run_pum.py \$home_dir 16 \$home_dir/results_attacks/pum
  • DMA: python3 \$home_dir/impact/covert_channel_attack/script/run_dma.py \$home_dir 16 \$home_dir/results_attacks/dma
A.5.5 Running Attacks for Figure 8 and 10 (PNM POC)
python3 \$home_dir/impact/covert_channel_attack/script/run_pnm_poc.py \$home_dir 16 \$home_dir/results_attacks/pnm_poc
A.5.6 Running Attacks for Figure 8 and 10 (PUM POC)
python3 \$home_dir/impact/covert_channel_attack/script/run_pum_poc.py \$home_dir 16 \$home_dir/results_attacks/pum_poc
A.5.7 Running Defense Mechanism Experiments for Figure 12
  • Download traces (if needed).
  • Create the jobfile: python3 \$home_dir/scripts/create_jobfile_impact.py \$home_dir/simulator/sniper \$home_dir/scripts/defenses.jobfile
  • Execute the jobfile: sh \$home_dir/scripts/defenses.jobfile
A.6 Plotting the Results
Once the experiments are completed, the reader can plot the figures using the plotting scripts we provide.
To reproduce the figures, 1) install plotting dependencies:
{mdframed}
[backgroundcolor=darkgray, linecolor=black, linewidth=0.5pt] \$ pip install numpy matplotlib pandas
2) run the plotting script: plot_figures.sh
{mdframed}
[backgroundcolor=darkgray, linecolor=black, linewidth=0.5pt] \$ sh plot_figures.sh
This script will create Figures 8, 9, 10, and 12 under the figures/ directory.
A.7 Evaluation & Expected Results
A.7.1 Post-Execution Steps
Review Figures 8, 9, 10, and 12 in the figures/ directories.
Note that the attack codes are compiled during the setup, and the results presented in Figures 8, 9, and 10 might vary slightly based on your system configurations (e.g., compiler version).
A.7.2 Troubleshooting
Refer to the troubleshooting section in the README.md for potential issues and solutions.
A.8 Methodology
Submission, reviewing and badging methodology: